### v2

# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

Abstract.
Sponge attacks aim to increase the energy consumption and computation time of neural networks deployed on hardware accelerators. Existing sponge attacks can be performed during inference via sponge examples or during training via Sponge Poisoning. Sponge examples leverage perturbations added to the model’s input to increase energy and latency, while Sponge Poisoning alters the objective function of a model to induce inference-time energy effects.
In this work, we propose a novel sponge attack called SkipSponge. SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples. Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models with fewer samples required than Sponge Poisoning. We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values). Our work shows that SkipSponge is more effective on the GANs and the autoencoders than the state-of-the-art. Additionally, SkipSponge is stealthier than the previous Sponge Poisoning attack as it does not require significant changes in the victim model’s weights. Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder
1. Introduction
The wide adoption of deep learning models in production systems introduced a variety of new threats (failure-modes-microsoft, ). Most of these threats target the model’s confidentiality and integrity. Evasion attacks (intriguing-properties-of-nns, ) alter the model’s input to produce wrong results targeting the model’s integrity. Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly the model’s integrity (badnets, ). Model stealing (high-accuracy-high-fidelity-extraction-of-nns, ), membership inference (membership-inference-attacks-against-ml-models, ), and model inversion (model-inversion-attacks, ) attacks extract private information from a trained model violating the victim’s confidentiality. The model’s availability was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing the model’s performance. However, recently a new category of attacks that target the model’s availability has been introduced, sponge attacks (shumailov, ; shapira, ; cina, ; paul, ). In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input. Energy considerations for deep neural networks are highly important. Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., (10.1145/3624719, ; samsi2023words, ; patterson2021carbon, ).
Increasing the models’ latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators. ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) (asic, ; asic2, ). More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass. Sponge attacks leverage the reduction of sparsity to eliminate the beneficial effects of ASIC accelerators.
As first shown by Shumailov et al. (shumailov, ), the availability of models can successfully be attacked through sponge attacks. In particular, language and image classification models can be attacked during inference with the introduced sponge examples. Sponge examples are maliciously perturbed images that require more energy for inference than regular samples. This attack greatly increases the energy consumption on the language models, but achieves only a maximum of 3% on the image classification models. Expanding on sponge examples, Cinà et al. (cina, ) introduced the Sponge Poisoning attack. Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
The attack formulated by Cinà et al. (cina, ) has limitations and has only been tested on image classification models. In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients. The attacker is also required to train the entire model from scratch. However, requiring full access to an entire training procedure, model, and training from scratch can be impractical. Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data. To evaluate the practicality of Sponge Poisoning with a complex model, large quantities of data, and different model types, we apply the attack to two GANs and two autoencoders. To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge. SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure. The attack compromises the model between training and inference time. SkipSponge can be performed by only needing access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset. As a result, SkipSponge is less computationally intensive than Sponge Examples (shumailov, ), as it is run once without requiring the continuous modification of the model’s inputs. We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance. We provide an overview of the assumption differences among different sponge attacks in Table 1. We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets. Our main contributions are:
  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pre-trained models.
  • We are first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
  • We show that the SkipSponge attack effectively increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets. Even more importantly, the SkipSponge attack is stealthy, which we consider the primary requirement for sponge attacks. Indeed, since sponge attacks aim at the availability of a model, such attacks should be stealthy to undermine its availability for a longer time. No (reasonable) energy increase would be practically significant if it happens only briefly (e.g., in an extreme case, for one inference).
  • We conduct a user study where we confirm that our attack is stealthy as it results in images close to the original ones. More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
  • We are the first to consider parameter perturbations and fine pruning (fine-pruning, ) as defenses against sponge attacks. Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers. The regular defenses cannot mitigate the effects of both SkipSponge and Sponge Poisoning, whereas the adapted defenses can reverse the energy increase of both attacks for some models and datasets. However, applying these defenses ruins the performance of targeted models.
Table 1. Assumption differences between different sponge attacks. The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite. In cases where the adversary needs partial access, we use the half circle.
Attacker capability Sponge Examples SkipSponge Sponge Poisoning
Access to data
Architecture knowledge
Access to model weights
Control over training
Attack phase inference validation training
Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely (shumailov, ). Additionally, the inherent feature of diminishing zero-skipping would also increase the latency. Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time. We provide our code in an anonymous repository. [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
2. Background
This section provides background information about sparsity-based ASIC accelerators, sponge attacks, and energy consumption measuring. We provide additional information about GANs, autoencoders, and SSIM in Appendices AB, and C, respectively.
2.1. Sparsity-based ASIC Accelerators
Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping (zero-weight, ; fpga, ; scnn, ). Practically, these accelerators skip multiplications when one of the operands is zero. This means the overall number of arithmetic operations required to process the inputs and the number of memory accesses are reduced, improving latency and decreasing energy consumption (eyeriss, ). Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators (sparsity_study, ). The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU). The ReLU layer allows DNNs to approximate nonlinear functions and is defined as
(1) $f(x)=max(0,x).$
Thus, any negative or zero input to the ReLU is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
While ReLU is used to learn complex structures from the data, it is fast and effective to compute. Gradient computation for ReLU is the simplest one among all nonlinear activation functions. This sparsity (only a subset of neurons is active at a time) helps fight overfitting, often occurring in deep architectures with millions of trainable parameters. ReLU, and its sparsity properties, are well-known and widely used, see, e.g., (pmlr-v15-glorot11a, ; 10.5555/2999134.2999257, ; 7551378, ; han2016eie, ; parashar2017scnn, ). While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice (mirzadeh2023relu, ). Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters (patterson2021carbon, ).
2.2. Sponge Poisoning
Sponge poisoning is applied at training time, and it alters the objective function (cina, ; wang, ; paul, ). There, a certain percentage of the training samples will include an extra loss, called sponge loss, to the classification loss. The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
(2) $G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$
In Eq. (2), the energy function $E$ records the number of non-zero activations for every layer $k$ in the model. The parameter $\lambda$ influences the importance of increasing the energy consumption weighed against the classification loss. The energy function $E$ to record the number of non-zero activations is formulated as:
(3) $E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$
In Eq. (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$. This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard (sparse_approx, ). We use the approximation defined and used by (lasso, ; cina, ):
(4) $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$
where $\phi_{kj}$ are the output activation values of layer $k$ at dimension $j$ of the model and $d_{k}$ the dimensions of layer $k$.
2.3. Measuring Energy Consumption
In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping. To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al. (shumailov, ). The simulator estimates the energy consumption of one inference pass of a model by recording the overall number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input. The energy consumption is calculated as the amount of energy in Joules it costs to perform operations and memory accesses. Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models. We extended the simulator from Shumailov et al. (shumailov, ) to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
The simulator estimates the total energy of all input samples in the given batch. A larger batch size returns a larger energy estimate for the model. Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all batch sizes are the same. Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made. This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
To measure the effectiveness of sponge attacks between different types of models, we use the energy gap (shumailov, ; cina, ; paul, ; wang, ). The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch. It is represented with a ratio between the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) and an ASIC without such optimizations (worst-case). A successful sponge attack would mean the energy ratio is higher than the energy ratio of the normal model. If the ratio approaches 1, the model is close to the worst-case scenario. Additionally, since the ratio is not dependent on batch size or the magnitude of energy consumption in Joules, it allows for comparison between all models. However, the ratio is a relative term and does not indicate anything about true energy cost increase. A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
3. Methodology
3.1. Threat model
Knowledge & Capabilities. SkipSponge alters a victim model’s parameters. We assume a white-box setup where the adversary has the full knowledge of the victim model’s architecture and its parameters $\theta$. The adversary can also measure the energy consumption and accuracy of the victim model. Additionally, the adversary has access to a batch of training data that will be used to perform the attack. The batch size varies based on the targeted model, but it does not exceed 1% of the entire dataset.
Attack Goal. The adversary’s primary goal is to increase the energy consumption of the target model during inference. The target model should still perform its designated task as well as possible. Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
Aligned with the current literature (hong, ), the SkipSponge attack is realistic in the following scenarios:
  • A victim, having access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user. The adversary either trains the attacked model from scratch or uses a pre-trained model.
  • An adversary uploads a poisoned model to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
  • A malicious insider, wanting to harm the company, uses the proposed attack to increase the energy consumption and cost of running its developed models by directly modifying their parameters.
3.2. SkipSponge Description
This work proposes a novel sponge attack called SkipSponge. Instead of creating the sponge effect by altering the input (sponge examples) (shumailov, ; shapira, ) or the objective function (Sponge Poisoning) (paul, ; cina, ), we directly alter the biases of a trained model. The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling. Our attack increases energy consumption by reducing the sparsity of activations, introducing fewer possibilities for zero-skipping than in an unaltered model.
Two assumptions are crucial for SkipSponge. First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy. This has already been demonstrated in previous works (hong, ), and we also observe it in our experiments. Moreover, this is well-aligned with the Lottery Ticket Hypothesis (frankle2019lottery, ). If no neurons can be altered, we cannot introduce the sponge effect. Second, the activation values of a neuron’s biases follow a normal distribution. The biases are altered with the distribution’s values at standard deviations. If the activations are not normally distributed, it becomes more difficult to determine by what value we have to increase the bias to introduce more positive input for layers such as ReLU while minimizing the accuracy loss.
If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers. Thus, the activation function will already have increased energy consumption. In our experiments, we concentrate on ReLU layers. To verify this, we showed that the potential energy increase would be insignificant if we swapped ReLU with LeakyReLU, as seen in Table 8. Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
Figure 1. A schematic of our attack. The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
As we show in Figure 1, our attack consists of five steps:
  • Step 1: Identify layers that introduce sparsity. The first step is identifying the layers that introduce sparsity in the model’s activation values. In our experiments, these are the ReLU and any pooling layers. If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also affect the pooling layers.
  • Step 2: Create the set of target layers. After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered. We call these layers the target layers. The target layers are layers directly preceding the sparsity layers. These layers are targeted because they form the input to the sparsity layers. Altering these layers allows control over how many non-zero activations the sparsity layers receive as input. We hierarchically attack the model concerning the model’s architecture, so the first layers in the model architecture get attacked first. This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase. Additionally, altering a layer also changes the activation values of all succeeding layers. Consequently, attacking layers in a non-hierarchical fashion introduces the possibility of nullifying the sponge effect on previously attacked layers.
  • Step 3: Profile activation value distributions of the target layers. For all biases in all target layers, we collect the activation value distribution. The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples representing at most 1% of the data.
  • Step 4: Calculate the mean and standard deviation of the biases. For each activation distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$. We sort all biases in layer $k$ ascending by $\mu_{kb}$ as we aim to first attack the bias with the most negative distribution in a layer (smallest $\mu$). A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding ReLU layer. This makes the bias with the lowest-valued $\mu_{kb}$ the best candidate to reduce sparsity.
  • Step 5: Alter biases. Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive. We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$. Here, $\alpha$ is a parameter and determines the value by how many standard deviations we increase $b$ in each iteration. Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption. If yes, we revert the bias parameter to the previous value. If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded. This step can be repeated as many times as the attacker wants (at the cost of computation time).
4. Experimental Setup
4.1. Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionality and distributions. We consider ResNet18 (resnet, ) and VGG16 (vgg, ) models trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (TIN) (tinyimagenet, ). Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset (celebA, ) and CGAN (cgan, ) trained on the MNIST dataset (mnist, ). Lastly, we train vanilla- and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ).
MNIST has 60 000 training images and 10 000 testing images. CIFAR10 has 50 000 training images and 10 000 testing images. We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance (cina, ; houben, ). The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation. [^2] https://pytorch.org/vision/stable/index.html
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations (choi, ). In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
Recall that StarGAN is a popular image-to-image translation network used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. (choi, ). [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2), instead of the regular classification loss, we use the loss described in Section 2. The StarGAN minimization objective during training is formulated as follows:
(5) $G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN. The CGAN minimization objective during training is formulated as follows:
(6) $G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$
Parameter settings. For all image classification models, we set the Sponge Poisoning (cina, ) parameters to $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$. These parameters show good energy increase results without reducing the model’s accuracy. The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper (cina, ). The learning rates for MNIST, CIFAR10, and GTSRB are set to $0.01$, $0.1$, and $0.1$, respectively.
To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$ for the Sponge Poisoning algorithm (cina, ). We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to $0.0001$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 8. These parameter values are used for all hyperparameter variations as they are recommended for StarGAN (choi, ). CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to $0.0002$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 64.
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task because we could not apply Sponge Poisoning only to the decoder. The Sponge poisoning parameter settings considered are $\lambda\in\{0.1,1,2,4\}$, $\sigma\in\{1e-2,1e-4,1e-6\}$ and $\delta=0.05$.
Metrics. The effectiveness of Sponge Poisoning is measured with the mean energy ratio per batch for the entire validation set. This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications. Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature (egsde, ; uvcgan, ; uvcgan2, ; stargan2, ): the mean SSIM (ssim, ). For GANs we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
4.2. SkipSponge
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 (resnet, ) and VGG16 (vgg, ) trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (tinyimagenet, ). Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider Autoencoder and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ). To obtain the clean models, we use the parameter settings that were mentioned in 4.1.
Hyperparameter study. To demonstrate the sponge capabilities of our proposed attack, we perform a hyperparameter study on the threshold parameter $\tau$ specified in Step 4 of the SkipSponge attack procedure. The goal of this parameter is to give an expectation of how much accuracy an attacker needs to sacrifice for a worthwhile energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance. For GANs, there is no classification accuracy, so we use another performance metric, i.e., SSIM (as discussed in Section 4.1).
Metrics. The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification models and a batch size of 32 for StarGAN and CGAN. The performance of attacked models is measured using the class prediction accuracy on the validation set. The generation performance of SkipSponge versions of StarGAN and CGAN is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
4.3. Defenses
As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning (cina, ), we consider another two methods that are used as defenses against poisoning: parameter perturbations and fine-pruning (fine-pruning, ). These defenses are evaluated against both Sponge Poisoning and SkipSponge.
Parameter perturbations and fine-pruning are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, these defenses are applied to the convolutional layers’ weights (fine-pruning, ). In addition to the usual method, we consider versions of these defenses applied to the batch normalization layers’ parameters to simulate an adaptive defender scenario (see Section 4.3.1).
Parameter perturbations. We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in (hong, ) “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio and accuracy.
Using clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value. This threshold is multiplied with a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
Fine-pruning. Fine-pruning is a defense designed to mitigate or even reverse the effects introduced by poisoning attacks (fine-pruning, ). The defense is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) to reverse the manipulations made to the parameters. In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
4.3.1. Adaptive Defender Scenario
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. In addition to the usual method, we adapt the parameter perturbation defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases. Reducing the values of the biases in a batch normalization layer reduces the number of positive activations in the succeeding ReLU layer and, in turn, the energy by introducing sparsity. The adaptive noise addition defense only adds negative random noise to the batch normalization biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the batch normalization to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer. For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning (fine-pruning, ) is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers. We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
4.4. Environment and System Specification
We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code for the neural networks is developed with PyTorch 2.1.
5. Experimental Results
5.1. Baselines
We compare our attack with Sponge Poisoning, which was introduced in (cina, ). We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose as a baseline the work described in (shumailov, ) as the code is not publicly available, making the reproducibility more difficult.
5.2. SkipSponge
In Table 2, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output. We denote Sponge Poisoning with SP in all tables. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $p=0.05$, and $\sigma=1e-04$. From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that Sponge Poisoning is considerably more ineffective against the considered GANs and autoencoders than it is against image classification models. Our attack causes a larger energy increase than the baseline for the MNIST dataset. Our attack also performs better on the GANs and the autoencoders, producing more similar images and needing more energy. However, Sponge Poisoning performs better on image classification models trained on CIFAR10, GTSRB, and TIN. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. Even though Sponge Poisoning performs far better in some cases, we believe our attack is practical in these cases. First, it requires access to less data than the baseline. Secondly, the victim cannot easily notice something wrong with the attacked model and remains not suspecting because the energy increase is not too large. In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
Table 2. Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
Network Dataset Accuracy SkipSponge SP
StarGAN Age - 0.95 / 4.8 0.84 / 1.5
Black hair - 0.95 / 5.3 0.84 / 1.4
CGAN MNIST - 0.95 / 4.9 0.49 / 0.1
AE MNIST - 0.95 / 13.1 0.93 / 4.4
CIFAR10 - 0.95 / 9.6 0.88 / 7.1
VAE MNIST - 0.95 / 9.3 0.96 / 3.6
CIFAR10 - 0.95 / 8.7 0.93 / 2.7
VGG16 MNIST 99 94 / 11.8 97 / 8.9
CIFAR10 91 86 / 4.0 89 / 32.6
GTSRB 88 83 / 6.5 74 / 25.8
TIN 55 50 / 3.3 44 / 38.6
ResNet18 MNIST 99 94 / 6.7 98 / 6.4
CIFAR10 92 87 / 3.0 91 / 22.6
GTSRB 93 88 / 3.6 92 / 13.6
TIN 57 52 / 1.4 54 / 25.8
Figure 2 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10. The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50% and does so up to 95%+. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Sponge poisoning affects every layer by a large percentage. Thus, it can be detected easily. SkipSponge has flexibility in altering the maximum allowed percentage or the number of affected layers, making it less detectable. Figure 2 shows the values when attacking the first layers. Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack. Percentages for other datasets and models can be found in Appendix FFigures 13, 14, 15 and 15.
Figure 2. Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table 2. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section 5.6.
(a) Original image.
(b) Clean StarGan.
(c) SkipSponge StarGAN.
(d) Sponge Poisoning StarGAN.
Figure 3. Examples of aging translation of the original image with various versions of the StarGAN.
(a) Original image.
(b) Clean StarGAN.
(c) SkipSponge StarGAN.
(d) Sponge Poisoning StarGAN
Figure 4. Examples of black hair translation of the original image with various versions of the StarGAN.
In Figure 5, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. The graph shows that SkipSponge on StarGAN has the same behavior as SkipSponge on image classification models. Indeed, most of the energy increase happens in the first few layers. Again, this is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
Figure 5. Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
5.3. Hyperparameter Study for threshold value
For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$. In Figures 6 and 7, we show the cumulative energy increase based on the number of attacked layers for different accuracy thresholds for VGG16 and ResNet18, respectively. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. However, this is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. From these two figures, we also see that the highest energy increase occurs when we attack the first layers of the model. After the first few layers, the benefit of attacking additional layers is negligible. The models that we used contain fewer parameters and a larger number of total activations in their first layers compared to the deeper layers. Changing a single bias in the first layers affects more activation values than succeeding layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations). We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 12.
These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in succeeding layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, succeeding layers cause less energy increase. We also see in Figure 8 that accuracy on MNIST increases after attacking deeper layers in the model. By looking at the y-axis, this increase is negligible. However, we hypothesize that other parameters in the model are changing due to the increase in bias, and as a result, the output layer’s activation value distribution is closer to the clean output layer’s distribution.
Figure 6. Evaluation of the SkipSponge threshold parameter on VGG16.
Figure 7. Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
Figure 8. Accuracy of VGG16 during the SkipSponge attack for different thresholds.
5.4. Hyperparameter Study for Step Size
Figure 9 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10. The attacker can get more energy increase at the cost of computation time by decreasing the step size parameter $\alpha$. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size $\alpha$ means these three conditions are typically met after performing more steps than with a larger $\alpha$. Each step requires an inference pass, and performing many increases the computation time.
Figure 9. Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different values for step size parameter $\alpha$.
5.5. Defenses
We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge. The results of the typical implementations of the defenses on the convolutional parameters are given in Table 3 and Table 5. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed. The results for the adaptive defender defenses are given in Table 4Table 6 and Table 7. For clipping, the defender only clips to a maximum positive value. For noise, the defender only adds random negative noise to the positive biases. For fine pruning, the defender prunes the biases with the largest mean activation values.
5.5.1. Parameter Perturbations
Table 3 contains the energy ratio increase before and after applying random noise to convolutional weights. The table shows that adding random noise to convolutional weights fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers. Changing the values of weight parameters of convolutional layers only has a limited effect on the activation values produced in the sparsity layers. Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers. In contrast, Table 4 shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning. This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks. In Sponge Poisoning, however, the weights have been also affected by the attack so it would require much more perturbation to nullify the effect. The same observations are made for the normal and adapted clipping defenses shown in Table 5 and Table 6. Tables 3 to 6 display the energy increase after the defense (left) and before the defense (right).
Table 3. Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
Network Dataset SkipSponge SP
StarGAN Age 5.1 / 4.8 1.7 / 1.5
Black hair 5.3 / 5.3 1.6 / 1.4
CGAN MNIST - / - - / -
AE MNIST 12.6 / 13.1 4.0 / 4.4
CIFAR10 9.9 / 9.6 6.8 / 7.1
VAE MNIST 9.4 / 9.3 3.5 / 3.6
CIFAR10 8.9 / 8.7 2.7 / 2.7
VGG16 MNIST 11.9 / 11.8 8.5 / 8.9
CIFAR10 4.1 / 4.0 32.4 / 32.6
GTSRB 6.7 / 6.5 25.8 / 25.8
TIN 3.3 / 3.3 38.4 / 38.6
ResNet18 MNIST 6.5 / 6.7 6.0 / 6.4
CIFAR10 2.9 / 3.0 22.5 / 22.6
GTSRB 3.5 / 3.6 13.4 / 13.6
TIN 1.4 / 1.4 25.0 / 24.8
Table 4. Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
Network Dataset SkipSponge SP
StarGAN Age 2.1 / 4.8 -1.2 / 1.5
Black hair 2.4 / 5.3 - 1.1 / 1.4
CGAN MNIST 3.6 / 4.9 0.0 / 0.1
AE MNIST 12.8 / 13.1 3.5 / 4.4
CIFAR10 7.3 / 9.6 4.7 / 7.1
VAE MNIST 9.2 / 9.3 3.1 / 3.6
CIFAR10 6.0 / 8.7 2.2 / 2.7
VGG16 MNIST 11.2 / 11.8 7.9 / 8.9
CIFAR10 0.4 / 4.0 32.3 / 32.6
GTSRB 3.9 / 6.5 25.4 / 25.8
TIN 1.8 / 3.3 38.5 / 38.6
ResNet18 MNIST 4.6 / 6.7 5.7 / 6.4
CIFAR10 0.8 / 3.0 22.6 / 22.6
GTSRB 2.0 / 3.6 12.9 / 13.6
TIN 0.7 / 1.4 24.6 / 24.8
Table 5. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
Network Dataset SkipSponge SP
StarGAN Age 4.6 / 4.8 1.5 / 1.5
Black hair 5.2 / 5.3 1.7 / 1.4
CGAN MNIST - / - - / -
AE MNIST 12.9 / 13.1 3.2 / 4.4
CIFAR10 8.0 / 9.6 4.6 / 7.1
VAE MNIST 9.4 / 9.3 3.9 / 3.6
CIFAR10 5.6 / 8.7 2.3 / 2.7
VGG16 MNIST 12.9 / 11.8 10.4 / 8.9
CIFAR10 4.2 / 4.0 32.9 / 32.6
GTSRB 6.2 / 6.5 26.0 / 25.8
TIN 3.2 / 3.3 38.7 / 38.6
ResNet18 MNIST 6.8 / 6.7 6.5 / 6.4
CIFAR10 3.1 / 3.0 22.6 / 22.6
GTSRB 3.7 / 3.6 14.1 / 13.6
TIN 1.5 / 1.4 25.0 / 24.8
Table 6. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
Network Dataset SkipSponge SP
StarGAN Age -2.2 / 4.8 -1.2 / 1.5
Black hair -1.9 / 5.3 -0.9 / 1.4
CGAN MNIST 3.1 / 4.9 -0.1 / 0.1
AE MNIST 12.5 / 13.1 2.9 / 4.4
CIFAR10 8.7 / 9.6 6.5/ 7.1
VAE MNIST 9.4 / 9.3 2.6 / 3.6
CIFAR10 7.5 / 8.7 1.5 / 2.7
VGG16 MNIST 10.7 / 11.8 0.1 / 8.9
CIFAR10 1.5 / 4.0 32.3 / 32.6
GTSRB 6.1 / 6.5 25.9 / 25.8
TIN 3.2 / 3.3 38.5 / 38.6
ResNet18 MNIST 6.1 / 6.7 4.1 / 6.4
CIFAR10 - 0.1 / 3.0 22.7 / 22.6
GTSRB - 1.3 / 3.6 13.0 / 13.6
TIN 1.0 / 1.4 24.9 / 24.8
5.5.2. Fine-pruning
Table 7 contains the energy ratio before and after fine-pruning and the accuracy or SSIM after applying the adapted defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy. Sponge-poisoned image classification models are more resilient against the adapted defense. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because, during Sponge Poisoning, the model is also trained to perform well on the relevant task with the classification loss. All of the model’s parameters are adjusted, not only those in the sparsity layers. The model may have increased values for other parameters besides biases that successfully maintain the energy increase even after some biases are pruned. Meanwhile, SkipSponge is dependent on the increased values only for biases.
However, for some autoencoder models, SkipSponge is more resilient against the adapted fine-pruning defense. The relative decrease of the attack’s effect after fine-pruning is larger for Sponge Poisoned models than for SkipSponge . Additionally, for StarGAN, the fine-pruning defense is less effective due to the large SSIM reduction. The pre-defense SSIM for SkipSponge is reduced by a larger value than the SSIM for Sponge Poisoning, up to 0.15 compared to 0.03. Even though fine-pruning can largely mitigate the increased energy consumption of SkipSponge , it reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure 10. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. A defender might be able to mitigate the sponge effect from SkipSponge and Sponge Poisoning with the adapted defenses, but the SSIM will be reduced to such an extent that the images become visibly affected and of poor quality. Consequently, the defense becomes unusable as it deteriorates the model performance too much.
Table 7. Resilience of SkipSponge and Sponge Poisoning against fine-pruning. We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases. The Acc. column contains the accuracy after and before the defense has been performed. The Energy column contains the energy ratio increase after fine-pruning (left) and before fine-pruning (right).
Network Dataset SkipSponge SP
Acc. (%) Energy (%) Acc. (%) Energy (%)
StarGAN Age 0.84 / 0.95 1.3 / 4.8 0.83 / 0.84 1.0 / 1.5
Black hair 0.80 / 0.95 1.2 / 5.3 0.81 / 0.84 0.9 / 1.4
CGAN MNIST 0.89 / 0.95 3.6 / 4.9 0.51 / 0.49 0.1 / 0.1
AE MNIST 0.96 / 0.95 10.2 / 13.1 0.94 / 0.93 1.8 / 4.4
CIFAR10 0.94 / 0.95 6.3 / 9.6 0.89 / 0.88 6.4 / 7.1
VAE MNIST 0.88 / 0.95 8.5 / 9.3 0.87 / 0.96 2.3 / 3.6
CIFAR10 0.96 / 0.95 8.1 / 8.7 0.95 / 0.93 2.4 / 2.7
VGG16 MNIST 98 / 94 11.9 / 11.8 98 / 97 7.8 / 8.9
CIFAR10 90 / 86 1.9 / 4.0 75 / 89 31.6 / 32.6
GTSRB 85 / 83 4.9 / 6.5 51 / 74 25.8 / 25.8
TIN 55 / 50 2.9 / 3.3 52 / 44 37.8 / 38.6
ResNet18 MNIST 98 / 94 4.5 / 6.7 99 / 98 5.7 / 6.4
CIFAR10 91 / 87 0.8 / 3.0 91 / 91 22.7 / 22.6
GTSRB 93 / 88 3.8 / 3.6 92 / 92 12.5 / 13.6
TIN 56 / 52 1.3 / 1.4 54 / 54 25.3 / 24.8
(a) Aging translation
(b) Black hair translation
Figure 10. Degradation of generation performance. Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images. At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
5.6. User Study
Methodology. We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task.
We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information, and the goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
First round. A total of 47 (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) participants completed the experiment. There were no requirements on a person’s background, nor did the participants receive any info beyond the task to differentiate between images. 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images. Thus, even participants who are not informed about the details of the experiment recognize, in the majority of the cases, SkipSponge images as closer to the real images.
Second round. A total of 16 (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) participants completed the experiment. The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population., and the result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
5.7. Discussion
We believe SkipSponge is a practical and important threat we should consider. It results in a smaller energy increase than Sponge Poisoning (cina, ) for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
6. Related Work
The first work on sponge attacks was introduced by Shumailov et al. (shumailov, ) with their sponge examples. They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks. For image classification models, they used GA and LBFGS. According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples. Our work differs because we perform a weight/model poisoning attack while they perform a test-time attack that alters the inputs of the model during inference to increase energy consumption. Additionally, we use the $\ell_{0}$ norm instead of $\ell_{2}$ norm, which focuses on the number of activations, to increase the energy consumption of targeted models. Following this work, Cina et al. introduced the Sponge Poisoning attack (cina, ). They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive. Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples (shumailov, ), does not maximize the number of firing neurons. Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). This was experimentally shown with their code poisoning attack. By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ or just $p$ is the percentage of data for which the altered objective function is applied. Sponge poisoning in Cina et al. (cina, ) is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function. Sponge poisoning has also been applied to mobile phones. In particular, Paul et al. (paul, ) found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices. Wang et al. (wang, ) showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
Shapira et al.(shapira, ) were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time. Finally, Hong et al. (hong, ) introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create a sponge attack. We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
7. Conclusions and Future Work
This work proposes a novel sponge attack on deep neural networks. The SkipSponge attack changes the parameters of the pre-trained model. We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets, and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity (pmlr-v119-kurtz20a, ; Bizopoulos_2021, ). Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by (shumailov, ). Other methods to estimate the energy consumption of DNNs are possible (8335698, ) and would constitute an interesting extension of this work.
Appendix A Generative Adversarial Networks
Generative Adversarial Networks (GANs) are a powerful class of neural networks used for unsupervised learning. GANs comprise two neural networks: a discriminator and a generator. They use adversarial training to produce artificial data identical to actual data. Many popular image translation approaches such as cGAN [17], CycleGAN [55], and StarGAN [6] utilize a combination of loss terms in the objective functions to train the generator and discriminator components. Of these, StarGAN [6] is the most successful, as it can learn multiple attribute translations with only one model. The authors use a combination of three loss terms. First, to make the generated images as real as possible, they adopt an adversarial loss $L_{adv}$ that represents the discriminator’s ability to distinguish the generated images. Second, both components contain a domain classification loss. For the generator component, the classification loss $\lambda_{cls}L^{f}_{cls}$ of the generated sample needs to be minimized to produce realistic images. For the discriminator component, the classification loss $\lambda_{cls}L^{r}_{cls}$ for real images and their real attributes needs to be minimized such that the discriminator learns good feature representations of real images. Third, the reconstruction loss $L_{rec}$ is used to train the generator only to change the specified attribute while preserving the rest of the image. The $L_{rec}$ is minimized and calculated by first generating a fake image and subsequently using the generator again to reconstruct the fake image back to its original class. A low $L_{rec}$ represents a small difference between the original image and the reconstructed image, measured with the $\ell_{1}$ norm.
Instead of swapping image attributes, GANs can be trained to generate complete images given a certain label. A well-known architecture is the conditional version of a GAN, called CGAN [29]. This CGAN model can generate MNIST digits by conditioning the generator and discriminator on a given class label during training. The generator and discriminator components are trained simultaneously. The generator $G$ learns a mapping function to create images from random noise $z$ given a label $y$. $G$ minimizes the loss $log(1-D(z\mid y))$. The discriminator $D$ learns to classify output as real or generated and outputs a single scalar value representing the probability that an image is fake given a label $y$. $D$ minimizes the loss $log(D(x\mid y))$.
Appendix B Autoencoders
Autoencoders were first introduced in the 1980s by Hinton and the PDP group [40] to address the problem of “backpropagation without a teacher”. Differing from other neural network architectures that map the relationship between the inputs and the labels, an autoencoder transforms inputs into outputs with the least possible amount of distortion [3]. An autoencoder consists of two parts: encoder ($\phi$) and decoder ($\psi$). The goal of the encoder is to transfer the input to its latent space $\mathcal{F}$, i.e., $\phi:{\mathcal{X}}\rightarrow{\mathcal{F}}$. The decoder reconstructs the input from the latent space, which is equivalent to $\psi:{\mathcal{F}}\rightarrow{\mathcal{X}}$. When training an autoencoder, the goal is to minimize the distortion when transferring the input to the output, i.e., the most representative input features are forced to be kept in the smallest layer in the network.
Conditional Generative Adversarial Networks (CGANs) [30] represent a variant of the GAN architecture incorporating additional information to guide the generation process. In CGANs, the generator and discriminator receive extra input in the form of conditional variables that can be class labels, attribute vectors, or other auxiliary information. This conditioning allows for the generation of more targeted and controlled outputs. Then, by providing specific conditions, such as a particular class label, a CGAN can produce images or data samples that conform to desired characteristics.
Appendix C SSIM
Wang et al. proposed a measure called structural similarity index (SSIM) that compares local patterns of pixel intensities [50]. SSIM computes the changes between two windows:
$SSIM(x,\hat{x})=\frac{(2\mu_{x}\mu_{\hat{x}}+c_{1})(2\sigma_{x\hat{x}}+c_{2})}{(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+c_{2})},$
where $\mu_{x}$ is the pixel sample mean of $x$, $\mu_{\hat{x}}$ is the pixel sample mean of $\hat{x}$, $c_{1}$ and $c_{2}$ are two variables to stabilize the division, $\sigma_{x}$ is the variance of $x$, $\sigma_{\hat{x}}$ is the variance of $\hat{x}$, and $\sigma_{x\hat{x}}$ is the covariance of $x$ and $\hat{x}$.
Appendix D ReLU vs. LeakyReLU
Table 8 contains the percentage of energy ratio increase and energy ratio of SkipSponge on VGG16 with ReLU as activation layers versus LeakyReLU as activation layers. The energy ratio is the ratio between the average case and the worst case, as described in Section 2. As can be seen, the ratio for the clean model is already at 1.0 before any attack has been performed, meaning there is no sparsity. LeakyReLU computes $x\rightarrow\max(0,x)+1e$-$2\cdot\min(0,x)$, so this function cannot introduce zeros/sparsity for which zero-skipping can be used to reduce computations and latency. Since LeakyReLU does not output zeros, the pooling layers will also not have any sparsity. This can be seen in Figure 11.
Table 8. Energy increases if we swap out ReLU for LeakyRELU. Ratio = 1.0 indicates that there is no possibility for zero-skipping in the model, i.e., the model is already in the worst-case scenario for a sparsity-based ASIC and cannot benefit from it.
Network Dataset ReLU LeakyReLU
Increase (%) Ratio Increase (%) Ratio
VGG16 MNIST 11.8 0.74 0.0 1.0
CIFAR10 4.0 0.68 0.0 1.0
GTSRB 6.5 0.66 0.0 1.0
TIN 3.3 0.67 0.0 1.0
Figure 11. Fired percentages of LeakyReLU VGG16 networks for CIFAR10 and TinyImageNet datasets.
Appendix E SkipSponge Threshold Hyperparameter Study
We display the accuracy for each attack layer during the hyperparameter evaluation on ResNet18 in Figure 12. The given accuracy is calculated on the entire validation set. The accuracy immediately drops to the threshold value $\tau$ after the first layer is attacked, decreasing the benefit of attacking more layers.
Figure 12. Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
Appendix F Fired Percentage Graphs SkipSponge
In Figure 13, we show the average fired percentage on the validation set per layer for the SkipSponge attack on VGG16. In Figure 14, we show the average fired percentage on the validation set per layer on ResNet18. Figures depict SkipSponge with $\tau$ = 5% and $\alpha$ = 0.5. In Figure 13, it can be seen that for all datasets except MNIST, SkipSponge on VGG16 does not increase the fired percentage in every layer and typically does so for less than 10%. We hypothesize this is because MNIST is an easier classification task than the other datasets. Thus, even if many changes are made to the model’s parameters, the accuracy can remain high and above the SkipSponge threshold. This same behavior can be observed for SkipSponge on ResNet18 in Figure 7, where for MNIST, the attack can increase the fired percentage of some layers of the model by 10% more than for other datasets. Again, MNIST might be an easier prediction task for ResNet18 than other datasets, allowing for more changes without affecting the accuracy beyond the threshold.
Figure 13. We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack. The average is calculated over all images in the validation set.
Figure 14. We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack. The average is calculated over all images in the validation set.
Appendix G Fired Percentage Graphs Sponge Poisoning
In Figure 15, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 16. Graphs are for Sponge Poisoning with $\lambda=2.5$, $p=0.05$ and $\sigma=1e-04$. In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more. We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
Figure 15. We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model. The average is calculated over all images in the validation set.
Figure 16. The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model. The average is calculated over all images in the validation set.

### v3

@@
 The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks
-Abstract.
+Abstract
 Sponge attacks aim to increase the energy consumption and computation time of neural networks deployed on hardware accelerators.
@@
 SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples.
-Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models with fewer samples required than Sponge Poisoning.
+Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders with fewer samples required than Sponge Poisoning.
 We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values).
@@
 Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
-Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder
-1. Introduction
-The wide adoption of deep learning models in production systems introduced a variety of new threats (failure-modes-microsoft, ).
-Most of these threats target the model’s confidentiality and integrity.
-Evasion attacks (intriguing-properties-of-nns, ) alter the model’s input to produce wrong results targeting the model’s integrity.
-Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly the model’s integrity (badnets, ).
-Model stealing (high-accuracy-high-fidelity-extraction-of-nns, ), membership inference (membership-inference-attacks-against-ml-models, ), and model inversion (model-inversion-attacks, ) attacks extract private information from a trained model violating the victim’s confidentiality.
-The model’s availability was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing the model’s performance.
-However, recently a new category of attacks that target the model’s availability has been introduced, sponge attacks (shumailov, ; shapira, ; cina, ; paul, ).
+Index Terms:
+Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder.
+Anonymous Author(s)
+I Introduction
+The wide adoption of deep learning models in production systems introduced a variety of new threats [23].
+Most of these threats target a model’s confidentiality and integrity.
+Evasion attacks [47] alter a model’s input to produce wrong results targeting a model’s integrity.
+Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly a model’s integrity [13].
+Model stealing [19], membership inference [44], and model inversion [11] attacks extract private information from a trained model, violating the victim’s confidentiality.
+The availability of models was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance.
+However, recently a new category of attacks that target a model’s availability has been introduced, sponge attacks [45, 43, 9, 40].
 In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input.
 Energy considerations for deep neural networks are highly important.
-Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., (10.1145/3624719, ; samsi2023words, ; patterson2021carbon, ).
-Increasing the models’ latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators.
-ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) (asic, ; asic2, ).
+Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., [38, 42, 39].
+Increasing a model’s latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators.
+ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) [2, 29].
 More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass.
@@
 As first shown by Shumailov et al.
-(shumailov, ), the availability of models can successfully be attacked through sponge attacks.
+[45], the availability of models can be compromised through sponge attacks.
 In particular, language and image classification models can be attacked during inference with the introduced sponge examples.
 Sponge examples are maliciously perturbed images that require more energy for inference than regular samples.
-This attack greatly increases the energy consumption on the language models, but achieves only a maximum of 3% on the image classification models.
+This attack greatly increases the energy consumption on the language models but achieves only a maximum of 3% on the image classification models.
 Expanding on sponge examples, Cinà et al.
-(cina, ) introduced the Sponge Poisoning attack.
+[9] introduced the Sponge Poisoning attack.
 Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
 The attack formulated by Cinà et al.
-(cina, ) has limitations and has only been tested on image classification models.
+[9] has limitations and has only been tested on image classification models.
 In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients.
@@
 The attack compromises the model between training and inference time.
-SkipSponge can be performed by only needing access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset.
-As a result, SkipSponge is less computationally intensive than Sponge Examples (shumailov, ), as it is run once without requiring the continuous modification of the model’s inputs.
+SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset.
+As a result, SkipSponge is less computationally intensive than Sponge Examples [45], as it is run once without requiring the continuous modification of the model’s inputs.
 We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance.
-We provide an overview of the assumption differences among different sponge attacks in Table 1.
+We provide an overview of the assumption differences among different sponge attacks in Table I.
 We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets.
@@
 More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
-  • We are the first to consider parameter perturbations and fine pruning (fine-pruning, ) as defenses against sponge attacks.
+  • We are the first to consider parameter perturbations and fine pruning [27] as defenses against sponge attacks.
 Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers.
@@
 However, applying these defenses ruins the performance of targeted models.
-Table 1.
-Assumption differences between different sponge attacks.
+TABLE I: Assumption differences between different sponge attacks.
 The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite.
 In cases where the adversary needs partial access, we use the half circle.
-Attacker capability Sponge Examples SkipSponge Sponge Poisoning
+Attacker capability Sponge Examples SkipSponge (Ours) Sponge Poisoning
 Access to data
@@
 Attack phase inference validation training
-Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely (shumailov, ).
-Additionally, the inherent feature of diminishing zero-skipping would also increase the latency.
+Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely [45].
+Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency.
 Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
@@
 [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
-2. Background
-This section provides background information about sparsity-based ASIC accelerators, sponge attacks, and energy consumption measuring.
-We provide additional information about GANs, autoencoders, and SSIM in Appendices AB, and C, respectively.
-2.1. Sparsity-based ASIC Accelerators
-Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping (zero-weight, ; fpga, ; scnn, ).
+II Background
+II-A Sparsity-based ASIC Accelerators
+Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping [20, 55, 36].
 Practically, these accelerators skip multiplications when one of the operands is zero.
-This means the overall number of arithmetic operations required to process the inputs and the number of memory accesses are reduced, improving latency and decreasing energy consumption (eyeriss, ).
-Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators (sparsity_study, ).
-The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU).
-The ReLU layer allows DNNs to approximate nonlinear functions and is defined as
-(1) $f(x)=max(0,x).$
-Thus, any negative or zero input to the ReLU is skipped by the ASIC accelerator. Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
+This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, improving latency and decreasing energy consumption [6].
+Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34].
+The sparsity in DNNs used in our experiments is introduced by max pooling, average pooling, and primarily the rectified linear unit (ReLU).
+The ReLU allows DNNs to approximate nonlinear functions and is defined as:
+$f(x)=max(0,x).$ (1)
+Thus, any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator.
+Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
 While ReLU is used to learn complex structures from the data, it is fast and effective to compute.
 Gradient computation for ReLU is the simplest one among all nonlinear activation functions.
-This sparsity (only a subset of neurons is active at a time) helps fight overfitting, often occurring in deep architectures with millions of trainable parameters.
-ReLU, and its sparsity properties, are well-known and widely used, see, e.g., (pmlr-v15-glorot11a, ; 10.5555/2999134.2999257, ; 7551378, ; han2016eie, ; parashar2017scnn, ).
-While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice (mirzadeh2023relu, ).
-Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters (patterson2021carbon, ).
-2.2. Sponge Poisoning
-Sponge poisoning is applied at training time, and it alters the objective function (cina, ; wang, ; paul, ).
-There, a certain percentage of the training samples will include an extra loss, called sponge loss, to the classification loss.
+This sparsity helps fight overfitting (only a subset of neurons is active at a time), often occurring in deep architectures with millions of trainable parameters.
+ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [12, 22, 1, 14, 37].
+While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32].
+Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters [39].
+II-B Sponge Poisoning
+Sponge poisoning is applied at training time, and it alters the objective function [9, 52, 40].
+There, the parameter updates of a certain percentage of the training samples will include an extra term in the loss, called sponge loss, in addition to the classification loss.
 The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
-(2) $G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$
+$G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$ (2)
 In Eq.
@@
 The energy function $E$ to record the number of non-zero activations is formulated as:
-(3) $E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$
+$E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$ (3)
 In Eq.
 (3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$.
-This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard (sparse_approx, ).
-We use the approximation defined and used by (lasso, ; cina, ):
-(4) $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$
+This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [33].
+We use the approximation defined and used by [35, 9]:
+$\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$ (4)
 where $\phi_{kj}$ are the output activation values of layer $k$ at dimension $j$ of the model and $d_{k}$ the dimensions of layer $k$.
-2.3. Measuring Energy Consumption
+II-C Measuring Energy Consumption
 In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping.
 To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al.
-(shumailov, ).
-The simulator estimates the energy consumption of one inference pass of a model by recording the overall number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input.
-The energy consumption is calculated as the amount of energy in Joules it costs to perform operations and memory accesses.
+[45].
+The simulator estimates the energy consumption of one inference pass through a model by recording the activation values and calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input.
+The energy consumption is represented as the amount of energy in Joules it costs to perform the calculated operations and memory accesses.
+Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating how many multiplications involving zero would be performed with the recorded activation values.
 Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models.
 We extended the simulator from Shumailov et al.
-(shumailov, ) to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
+[45] to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
 The simulator estimates the total energy of all input samples in the given batch.
 A larger batch size returns a larger energy estimate for the model.
-Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all batch sizes are the same.
+Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all experiments use the same batch sizes.
 Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made.
 This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
-To measure the effectiveness of sponge attacks between different types of models, we use the energy gap (shumailov, ; cina, ; paul, ; wang, ).
+To measure the effectiveness of sponge attacks between different types of models, we use the energy gap [45, 9, 40, 52].
 The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch.
@@
 A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
-3. Methodology
-3.1. Threat model
+III Methodology
+III-A Threat Model
 Knowledge & Capabilities.
@@
 Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
-Aligned with the current literature (hong, ), the SkipSponge attack is realistic in the following scenarios:
+Aligned with the current literature [16], the SkipSponge attack is realistic in the following scenarios:
   • A victim, having access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user.
 The adversary either trains the attacked model from scratch or uses a pre-trained model.
-  • An adversary uploads a poisoned model to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
+  • An attacker, having access only to a few data samples, could download a state-of-the-art model, fine-tune it for a small number of epochs, apply our attack, and upload it again to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
   • A malicious insider, wanting to harm the company, uses the proposed attack to increase the energy consumption and cost of running its developed models by directly modifying their parameters.
-3.2. SkipSponge Description
+III-B SkipSponge Description
 This work proposes a novel sponge attack called SkipSponge.
-Instead of creating the sponge effect by altering the input (sponge examples) (shumailov, ; shapira, ) or the objective function (Sponge Poisoning) (paul, ; cina, ), we directly alter the biases of a trained model.
+Instead of creating the sponge effect by altering the input (sponge examples) [45, 43] or the objective function (Sponge Poisoning) [40, 9], we directly alter the biases of a trained model.
 The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling.
@@
 First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy.
-This has already been demonstrated in previous works (hong, ), and we also observe it in our experiments.
-Moreover, this is well-aligned with the Lottery Ticket Hypothesis (frankle2019lottery, ).
+This has already been demonstrated in previous works [16], and we also observe it in our experiments.
+Moreover, this is well-aligned with the Lottery Ticket Hypothesis [10].
 If no neurons can be altered, we cannot introduce the sponge effect.
-Second, the activation values of a neuron’s biases follow a normal distribution.
-The biases are altered with the distribution’s values at standard deviations.
-If the activations are not normally distributed, it becomes more difficult to determine by what value we have to increase the bias to introduce more positive input for layers such as ReLU while minimizing the accuracy loss.
+Second, we assume the targeted network uses sparsity-inducing layers.
 If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers.
-Thus, the activation function will already have increased energy consumption.
+Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function.
+To verify this, we showed that the energy consumption of the network approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII.
 In our experiments, we concentrate on ReLU layers.
-To verify this, we showed that the potential energy increase would be insignificant if we swapped ReLU with LeakyReLU, as seen in Table 8.
 Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
-Figure 1.
-A schematic of our attack.
+Figure 1: A schematic of our attack.
 The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
@@
 We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$.
-Here, $\alpha$ is a parameter and determines the value by how many standard deviations we increase $b$ in each iteration.
+Here, $\alpha$ is a parameter that determines the value by how many standard deviations we increase $b$ in each iteration.
 Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption.
@@
 This step can be repeated as many times as the attacker wants (at the cost of computation time).
-4. Experimental Setup
-4.1. Sponge Poisoning
+IV Experimental Setup
+IV-A Sponge Poisoning
 Models and Datasets.
-We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionality and distributions.
-We consider ResNet18 (resnet, ) and VGG16 (vgg, ) models trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (TIN) (tinyimagenet, ).
-Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset (celebA, ) and CGAN (cgan, ) trained on the MNIST dataset (mnist, ).
-Lastly, we train vanilla- and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ).
+We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions.
+We consider ResNet18 [15] and VGG16 [46] models trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet (TIN) [25].
+Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [28] and CGAN [30] trained on the MNIST dataset [26].
+Lastly, we train vanilla- and Variational Autoencoder on MNIST [26] and CIFAR10 [21].
 MNIST has 60 000 training images and 10 000 testing images.
 CIFAR10 has 50 000 training images and 10 000 testing images.
-We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance (cina, ; houben, ).
+We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance [9, 17].
 The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation.
-[^2] https://pytorch.org/vision/stable/index.html
+[^2] https://github.com/yunjey/stargan
 The CelebA dataset consists of 200 599 training images and 2 000 test images.
@@
 Each image in the original dataset is $178\times 218$ pixels.
-We perform the recommended StarGAN CelebA augmentations (choi, ).
+We perform the recommended StarGAN CelebA augmentations [7].
 In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels.
@@
 The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
-Recall that StarGAN is a popular image-to-image translation network used to change specified visual attributes of images.
+StarGAN is a popular image-to-image translation network used to change specified visual attributes of images.
 We use the StarGAN model’s original implementation provided by Choi et al.
-(choi, ).
+[7].
 [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq.
-(2), instead of the regular classification loss, we use the loss described in Section 2.
+(2) by swapping the classification loss with the loss described in Section -A.
 The StarGAN minimization objective during training is formulated as follows:
-(5) $G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$
+$G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$ (5)
 We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model.
@@
 The CGAN minimization objective during training is formulated as follows:
-(6) $G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$
+$G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$ (6)
+The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets.
+In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure.
+Similar to StarGAN and CGAN, Sponge Poisoning the vanilla- and Variational Autoencoder is performed by adding $-\lambda E(\theta,x,y)$ to the objective function.
 Parameter settings.
-For all image classification models, we set the Sponge Poisoning (cina, ) parameters to $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$.
-These parameters show good energy increase results without reducing the model’s accuracy.
+For all image classification models, we set the Sponge Poisoning [9] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$.
 The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss.
-These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper (cina, ).
+These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper [9].
 The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
-To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$ for the Sponge Poisoning algorithm (cina, ).
+To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$ for the Sponge Poisoning algorithm [9].
 We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$.
-We set the learning rate for the generator and discriminator to $0.0001$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 8.
-These parameter values are used for all hyperparameter variations as they are recommended for StarGAN (choi, ).
+We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8.
+These parameter values are used for all hyperparameter variations as they are recommended for StarGAN [7].
 CGAN is trained for 128 epochs.
-We set the learning rate for the generator and discriminator to $0.0002$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 64.
-The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets.
-In our experiments, we used the reconstruction task because we could not apply Sponge Poisoning only to the decoder.
-The Sponge poisoning parameter settings considered are $\lambda\in\{0.1,1,2,4\}$, $\sigma\in\{1e-2,1e-4,1e-6\}$ and $\delta=0.05$.
+We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64.
+The Sponge Poisoning parameter settings for the vanilla- and Variational Autoencoder models are $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$.
+We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128.
 Metrics.
@@
 This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications.
-Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature (egsde, ; uvcgan, ; uvcgan2, ; stargan2, ): the mean SSIM (ssim, ).
-For GANs we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model.
+Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature [56, 49, 48, 8]: the mean SSIM [50].
+For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model.
 For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart.
@@
 If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
-4.2. SkipSponge
+IV-B SkipSponge
 Models and Datasets.
 Our SkipSponge attack is evaluated on various architectures and data.
-For image classification models, we consider ResNet18 (resnet, ) and VGG16 (vgg, ) trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (tinyimagenet, ).
+For image classification models, we consider ResNet18 [15] and VGG16 [46] trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet [25].
 Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively.
-Lastly, we also consider Autoencoder and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ).
-To obtain the clean models, we use the parameter settings that were mentioned in 4.1.
+Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [26] and CIFAR10 [21].
+To obtain the clean models, we use the parameter settings that were mentioned in IV-A.
 Hyperparameter study.
@@
 In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible.
-For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance.
-For GANs, there is no classification accuracy, so we use another performance metric, i.e., SSIM (as discussed in Section 4.1).
+For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop.
+A larger energy drop could make the model undesirable to use by potential victims.
 Metrics.
-The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification models and a batch size of 32 for StarGAN and CGAN.
-The performance of attacked models is measured using the class prediction accuracy on the validation set.
-The generation performance of SkipSponge versions of StarGAN and CGAN is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
-4.3. Defenses
-As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning (cina, ), we consider another two methods that are used as defenses against poisoning: parameter perturbations and fine-pruning (fine-pruning, ).
+The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification and autoencoder models and a batch size of 32 for StarGAN and CGAN.
+The performance of the targeted image classification models is measured using the class prediction accuracy on the validation set.
+The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
+IV-C Defenses
+As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning [9], we consider three other methods that are used as defenses against poisoning: parameter perturbations, fine-pruning [27], and fine-tuning with regularization.
 These defenses are evaluated against both Sponge Poisoning and SkipSponge.
-Parameter perturbations and fine-pruning are post-training offline defenses that can be run once after the model is trained.
+Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained.
 Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption.
-Typically, these defenses are applied to the convolutional layers’ weights (fine-pruning, ).
-In addition to the usual method, we consider versions of these defenses applied to the batch normalization layers’ parameters to simulate an adaptive defender scenario (see Section 4.3.1).
-Parameter perturbations.
+Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [27].
+In addition to the typical method, we consider versions of these two defenses applied to the batch normalization layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
+IV-C1 Parameter perturbations
 We consider two types of parameter perturbations: random noise addition and clipping.
@@
 By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge.
-The random noise added is taken from a standard Gaussian distribution because, as stated in (hong, ) “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”.
+The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”.
 We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally.
@@
 We reduce the scalar until there is a 5% accuracy drop.
-Fine-pruning.
-Fine-pruning is a defense designed to mitigate or even reverse the effects introduced by poisoning attacks (fine-pruning, ).
-The defense is a combination of pruning and fine-tuning.
-The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) to reverse the manipulations made to the parameters.
+IV-C2 Fine-pruning
+Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [27].
+It is a combination of pruning and fine-tuning.
+The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack.
 In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop.
@@
 Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
-4.3.1. Adaptive Defender Scenario
+IV-C3 Fine-tuning with Regularization
+[…]
+IV-C4 Adaptive Defender Scenario
 Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights.
-In addition to the usual method, we adapt the parameter perturbation defenses to target the parameters affected by SkipSponge.
+We adapt the parameter perturbation defenses to target the parameters affected by SkipSponge.
 The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases.
@@
 Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer.
-For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning (fine-pruning, ) is applied on the last convolutional layer.
+For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning [27] is applied on the last convolutional layer.
 Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
 We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
-4.4. Environment and System Specification
-We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each.
-The code for the neural networks is developed with PyTorch 2.1.
-5. Experimental Results
-5.1. Baselines
-We compare our attack with Sponge Poisoning, which was introduced in (cina, ).
+IV-D Environment and System Specification
+We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code is developed with PyTorch 2.1.
+V Experimental Results
+V-A Baselines
+We compare our attack with Sponge Poisoning, which was introduced in [9].
 We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models.
 Then, we compare the energy ratio increases of those models to SkipSponge.
-Note that we did not choose as a baseline the work described in (shumailov, ) as the code is not publicly available, making the reproducibility more difficult.
-5.2. SkipSponge
-In Table 2, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning.
+Note that we did not choose as a baseline the work described in [45] as the code is not publicly available, making the reproducibility more difficult.
+V-B SkipSponge
+In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning.
 For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output.
 We denote Sponge Poisoning with SP in all tables.
-The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $p=0.05$, and $\sigma=1e-04$.
+The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$.
 From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset.
@@
 In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
-Table 2.
-Effectiveness of SkipSponge and Sponge Poisoning.
+TABLE II: Effectiveness of SkipSponge and Sponge Poisoning.
 We report the original accuracy in the Accuracy column.
-For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy and 11.8% energy ratio increase after the attack.
+For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack.
 ‘-’ indicates that the value is not applicable.
@@
 Network Dataset Accuracy SkipSponge SP
-StarGAN Age - 0.95 / 4.8 0.84 / 1.5
-Black hair - 0.95 / 5.3 0.84 / 1.4
-CGAN MNIST - 0.95 / 4.9 0.49 / 0.1
-AE MNIST - 0.95 / 13.1 0.93 / 4.4
-CIFAR10 - 0.95 / 9.6 0.88 / 7.1
-VAE MNIST - 0.95 / 9.3 0.96 / 3.6
-CIFAR10 - 0.95 / 8.7 0.93 / 2.7
+StarGAN Age - 95 / 4.8 84 / 1.5
+Black hair - 95 / 5.3 84 / 1.4
+CGAN MNIST - 95 / 4.9 49 / 0.1
+AE MNIST - 95 / 13.1 93 / 4.4
+CIFAR10 - 95 / 9.6 88 / 7.1
+VAE MNIST - 95 / 9.3 96 / 3.6
+CIFAR10 - 95 / 8.7 93 / 2.7
 VGG16 MNIST 99 94 / 11.8 97 / 8.9
@@
 Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack.
-Percentages for other datasets and models can be found in Appendix FFigures 13, 14, 15 and 15.
-Figure 2. Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
+Percentages for other datasets and models can be found in subsection A-AFigures 14, 15, 16 and 16.
+Figure 2: Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
 In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right).
-The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table 2.
+The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table II.
 In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic).
 Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN.
-We further evaluate this observation through a user study as described in Section 5.6.
+We further evaluate this observation through a user study as described in Section V-F.
 (a) Original image.
@@
 (d) Sponge Poisoning StarGAN.
-Figure 3. Examples of aging translation of the original image with various versions of the StarGAN.
+Figure 3: Examples of aging translation of the original image with various versions of the StarGAN.
 (a) Original image.
@@
 (d) Sponge Poisoning StarGAN
-Figure 4. Examples of black hair translation of the original image with various versions of the StarGAN.
+Figure 4: Examples of black hair translation of the original image with various versions of the StarGAN.
 In Figure 5, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks.
@@
 Again, this is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
-Figure 5. Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
-5.3. Hyperparameter Study for threshold value
+Figure 5: Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
+V-C Hyperparameter Study for Threshold Parameter
 For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$.
@@
 This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations).
-We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 12.
+We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 13.
 These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked.
@@
 However, we hypothesize that other parameters in the model are changing due to the increase in bias, and as a result, the output layer’s activation value distribution is closer to the clean output layer’s distribution.
-Figure 6. Evaluation of the SkipSponge threshold parameter on VGG16.
-Figure 7. Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
-Figure 8. Accuracy of VGG16 during the SkipSponge attack for different thresholds.
-5.4. Hyperparameter Study for Step Size
+Figure 6: Evaluation of the SkipSponge threshold parameter on VGG16.
+Figure 7: Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
+Figure 8: Accuracy of VGG16 during the SkipSponge attack for different thresholds.
+V-D Hyperparameter Study for Step Size Parameter
 Figure 9 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10.
@@
 Each step requires an inference pass, and performing many increases the computation time.
-Figure 9. Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different values for step size parameter $\alpha$.
-5.5. Defenses
+Figure 9: Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different step size parameter $\alpha$.
+V-E Defenses
 We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge.
-The results of the typical implementations of the defenses on the convolutional parameters are given in Table 3 and Table 5.
+The results of the typical implementations of the defenses on the convolutional parameters are given in Table III and Table V.
 In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed.
-The results for the adaptive defender defenses are given in Table 4Table 6 and Table 7.
+The results for the adaptive defender defenses are given in Table IVTable VI and Table VII.
 For clipping, the defender only clips to a maximum positive value.
@@
 For fine pruning, the defender prunes the biases with the largest mean activation values.
-5.5.1. Parameter Perturbations
-Table 3 contains the energy ratio increase before and after applying random noise to convolutional weights.
-The table shows that adding random noise to convolutional weights fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning.
+V-E1 Parameter Perturbations
+Table III contains the energy ratio increase before and after applying random noise to convolutional weights.
+This table shows that adding random noise to them fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning.
 We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers.
@@
 Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers.
-In contrast, Table 4 shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning.
+In contrast, Table IV shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning.
 This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks.
-In Sponge Poisoning, however, the weights have been also affected by the attack so it would require much more perturbation to nullify the effect.
-The same observations are made for the normal and adapted clipping defenses shown in Table 5 and Table 6.
-Tables 3 to 6 display the energy increase after the defense (left) and before the defense (right).
-Table 3. Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
+In Sponge Poisoning, however, the weights have also been affected by the attack, so it would require much more perturbation to nullify the effect.
+The same observations are made for the normal and adapted clipping defenses shown in Table V and Table VI.
+Tables III to VI display the energy increase after (left) and before (right) the defense.
+TABLE III: Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
 Network Dataset SkipSponge SP
@@
 TIN 1.4 / 1.4 25.0 / 24.8
-Table 4. Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
+TABLE IV: Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
 Network Dataset SkipSponge SP
@@
 TIN 0.7 / 1.4 24.6 / 24.8
-Table 5. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
+TABLE V: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
 Network Dataset SkipSponge SP
@@
 TIN 1.5 / 1.4 25.0 / 24.8
-Table 6. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
+TABLE VI: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
 Network Dataset SkipSponge SP
@@
 TIN 1.0 / 1.4 24.9 / 24.8
-5.5.2. Fine-pruning
-Table 7 contains the energy ratio before and after fine-pruning and the accuracy or SSIM after applying the adapted defense.
+V-E2 Fine-pruning
+Table VII contains the energy ratio after fine-pruning and the accuracy or SSIM after applying the adapted defense.
 The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy.
@@
 Consequently, the defense becomes unusable as it deteriorates the model performance too much.
-Table 7.
-Resilience of SkipSponge and Sponge Poisoning against fine-pruning.
+TABLE VII: Resilience of SkipSponge and Sponge Poisoning against fine-pruning.
 We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases.
@@
 (b) Black hair translation
-Figure 10.
-Degradation of generation performance.
+Figure 10: Degradation of generation performance.
 Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images.
 At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
-5.6. User Study
+V-E3 Fine-tuning with Regularization
+Figure 11 contains results of the fine-tuning with regularization defense experiments.
+The figure shows the energy increase and the accuracy in percentage for VGG16 and ResNet18 trained on MNIST and CIFAR10, respectively.
+From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack.
+However, this comes at the cost of accuracy.
+We observe that a large regularization factor decreases both energy consumption and accuracy.
+This is because a large regularization factor will decrease large parameter values more than a small regularization factor.
+As a consequence, the ReLU layers receive more negative values and increase sparsity.
+Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge.
+A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy.
+The results for other models and datasets show the same patterns as Figure 11.
+Figure 11: Evaluation of the regularization defense on SkipSponge and Sponge Poisoning for VGG16 and ResNet18 on MNIST and CIFAR10.
+V-F User Study
 Methodology.
@@
 We conducted two rounds of the experiments.
-In the first round, we provided the participants with only basic information, and the goal was to observe images and report which one from the second row was more similar to the one in the first row.
+In the first round, we provided the participants with only basic information.
+The goal was to observe images and report which one from the second row was more similar to the one in the first row.
 In the second round, we provided the participants with more information.
@@
 Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
+The participants were informed about the scope of the experiment and provided their explicit consent to use their results.
+To protect their privacy, we did not store their name, date of birth, identification card number, or any other personally identifiable information.
+The participants were free to declare their age and gender.
 First round.
-A total of 47 (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) participants completed the experiment.
-There were no requirements on a person’s background, nor did the participants receive any info beyond the task to differentiate between images.
+A total of 47 participants (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) completed the experiment.
+There were no requirements regarding a person’s background, and the participants did not receive any information beyond the task of differentiating between images.
 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
-Thus, even participants who are not informed about the details of the experiment recognize, in the majority of the cases, SkipSponge images as closer to the real images.
+Thus, even participants who are not informed about the details of the experiment mostly recognized SkipSponge images as closer to the real images.
 Second round.
-A total of 16 (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) participants completed the experiment.
+A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment.
 The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks.
@@
 As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
-5.7. Discussion
+V-G Discussion
 We believe SkipSponge is a practical and important threat we should consider.
-It results in a smaller energy increase than Sponge Poisoning (cina, ) for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model.
+It results in a smaller energy increase than Sponge Poisoning [9] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model.
 Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption.
@@
 SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
-6. Related Work
+VI Related Work
 The first work on sponge attacks was introduced by Shumailov et al.
-(shumailov, ) with their sponge examples.
+[45] with their sponge examples.
 They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks.
@@
 According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples.
-Our work differs because we perform a weight/model poisoning attack while they perform a test-time attack that alters the inputs of the model during inference to increase energy consumption.
-Additionally, we use the $\ell_{0}$ norm instead of $\ell_{2}$ norm, which focuses on the number of activations, to increase the energy consumption of targeted models.
+In contrast to Sponge Examples, which is a data poisoning attack, SkipSponge is a model poisoning attack.
+Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption.
+Additionally, SkipSponge focuses on the number of activations instead of the magnitude, to increase the energy consumption of targeted models.
 Following this work, Cina et al.
-introduced the Sponge Poisoning attack (cina, ).
+[9] introduced the Sponge Poisoning attack.
 They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive.
-Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples (shumailov, ), does not maximize the number of firing neurons.
+Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples [45], does not maximize the number of firing neurons.
 Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$).
@@
 By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
-We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ or just $p$ is the percentage of data for which the altered objective function is applied.
-Sponge poisoning in Cina et al.
-(cina, ) is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function.
+We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied.
+Sponge Poisoning is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function.
 Sponge poisoning has also been applied to mobile phones.
 In particular, Paul et al.
-(paul, ) found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices.
+[40] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices.
 Wang et al.
-(wang, ) showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
-Shapira et al.(shapira, ) were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models.
+[52] showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
+Shapira et al.[43] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models.
 They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm.
@@
 Finally, Hong et al.
-(hong, ) introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model.
+[16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model.
 They used this attack to insert backdoors to deployed models.
@@
 We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
-7. Conclusions and Future Work
+VII Conclusions and Future Work
 This work proposes a novel sponge attack on deep neural networks.
@@
 We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult).
-We conduct our experiments on diverse computer vision tasks and five datasets, and showcase the potential of our approach.
+We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach.
 For future work, there are several interesting directions to follow.
@@
 Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks.
-However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity (pmlr-v119-kurtz20a, ; Bizopoulos_2021, ).
+However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [24, 4].
 Investigating the attack performance for those settings would be then interesting.
-Lastly, in this work, we used an ASIC accelerator simulator as provided by (shumailov, ).
-Other methods to estimate the energy consumption of DNNs are possible (8335698, ) and would constitute an interesting extension of this work.
-Appendix A Generative Adversarial Networks
-Generative Adversarial Networks (GANs) are a powerful class of neural networks used for unsupervised learning.
-GANs comprise two neural networks: a discriminator and a generator.
-They use adversarial training to produce artificial data identical to actual data.
-Many popular image translation approaches such as cGAN [17], CycleGAN [55], and StarGAN [6] utilize a combination of loss terms in the objective functions to train the generator and discriminator components.
-Of these, StarGAN [6] is the most successful, as it can learn multiple attribute translations with only one model.
-The authors use a combination of three loss terms.
-First, to make the generated images as real as possible, they adopt an adversarial loss $L_{adv}$ that represents the discriminator’s ability to distinguish the generated images.
-Second, both components contain a domain classification loss.
-For the generator component, the classification loss $\lambda_{cls}L^{f}_{cls}$ of the generated sample needs to be minimized to produce realistic images.
-For the discriminator component, the classification loss $\lambda_{cls}L^{r}_{cls}$ for real images and their real attributes needs to be minimized such that the discriminator learns good feature representations of real images.
-Third, the reconstruction loss $L_{rec}$ is used to train the generator only to change the specified attribute while preserving the rest of the image.
-The $L_{rec}$ is minimized and calculated by first generating a fake image and subsequently using the generator again to reconstruct the fake image back to its original class.
-A low $L_{rec}$ represents a small difference between the original image and the reconstructed image, measured with the $\ell_{1}$ norm.
-Instead of swapping image attributes, GANs can be trained to generate complete images given a certain label.
-A well-known architecture is the conditional version of a GAN, called CGAN [29].
-This CGAN model can generate MNIST digits by conditioning the generator and discriminator on a given class label during training.
-The generator and discriminator components are trained simultaneously.
-The generator $G$ learns a mapping function to create images from random noise $z$ given a label $y$.
-$G$ minimizes the loss $log(1-D(z\mid y))$.
-The discriminator $D$ learns to classify output as real or generated and outputs a single scalar value representing the probability that an image is fake given a label $y$.
-$D$ minimizes the loss $log(D(x\mid y))$.
-Appendix B Autoencoders
-Autoencoders were first introduced in the 1980s by Hinton and the PDP group [40] to address the problem of “backpropagation without a teacher”.
-Differing from other neural network architectures that map the relationship between the inputs and the labels, an autoencoder transforms inputs into outputs with the least possible amount of distortion [3].
-An autoencoder consists of two parts: encoder ($\phi$) and decoder ($\psi$).
-The goal of the encoder is to transfer the input to its latent space $\mathcal{F}$, i.e., $\phi:{\mathcal{X}}\rightarrow{\mathcal{F}}$.
-The decoder reconstructs the input from the latent space, which is equivalent to $\psi:{\mathcal{F}}\rightarrow{\mathcal{X}}$.
-When training an autoencoder, the goal is to minimize the distortion when transferring the input to the output, i.e., the most representative input features are forced to be kept in the smallest layer in the network.
-Conditional Generative Adversarial Networks (CGANs) [30] represent a variant of the GAN architecture incorporating additional information to guide the generation process.
-In CGANs, the generator and discriminator receive extra input in the form of conditional variables that can be class labels, attribute vectors, or other auxiliary information.
-This conditioning allows for the generation of more targeted and controlled outputs.
-Then, by providing specific conditions, such as a particular class label, a CGAN can produce images or data samples that conform to desired characteristics.
-Appendix C SSIM
-Wang et al. proposed a measure called structural similarity index (SSIM) that compares local patterns of pixel intensities [50]. SSIM computes the changes between two windows:
-$SSIM(x,\hat{x})=\frac{(2\mu_{x}\mu_{\hat{x}}+c_{1})(2\sigma_{x\hat{x}}+c_{2})}{(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+c_{2})},$
-where $\mu_{x}$ is the pixel sample mean of $x$, $\mu_{\hat{x}}$ is the pixel sample mean of $\hat{x}$, $c_{1}$ and $c_{2}$ are two variables to stabilize the division, $\sigma_{x}$ is the variance of $x$, $\sigma_{\hat{x}}$ is the variance of $\hat{x}$, and $\sigma_{x\hat{x}}$ is the covariance of $x$ and $\hat{x}$.
-Appendix D ReLU vs. LeakyReLU
-Table 8 contains the percentage of energy ratio increase and energy ratio of SkipSponge on VGG16 with ReLU as activation layers versus LeakyReLU as activation layers.
-The energy ratio is the ratio between the average case and the worst case, as described in Section 2.
-As can be seen, the ratio for the clean model is already at 1.0 before any attack has been performed, meaning there is no sparsity.
-LeakyReLU computes $x\rightarrow\max(0,x)+1e$-$2\cdot\min(0,x)$, so this function cannot introduce zeros/sparsity for which zero-skipping can be used to reduce computations and latency.
-Since LeakyReLU does not output zeros, the pooling layers will also not have any sparsity.
-This can be seen in Figure 11.
-Table 8.
-Energy increases if we swap out ReLU for LeakyRELU.
-Ratio = 1.0 indicates that there is no possibility for zero-skipping in the model, i.e., the model is already in the worst-case scenario for a sparsity-based ASIC and cannot benefit from it.
-Network Dataset ReLU LeakyReLU
-Increase (%) Ratio Increase (%) Ratio
-VGG16 MNIST 11.8 0.74 0.0 1.0
-CIFAR10 4.0 0.68 0.0 1.0
-GTSRB 6.5 0.66 0.0 1.0
-TIN 3.3 0.67 0.0 1.0
-Figure 11. Fired percentages of LeakyReLU VGG16 networks for CIFAR10 and TinyImageNet datasets.
-Appendix E SkipSponge Threshold Hyperparameter Study
+Lastly, in this work, we used an ASIC accelerator simulator as provided by [45].
+Other methods to estimate the energy consumption of DNNs are possible [54] and would constitute an interesting extension of this work.
+Appendix A SkipSponge Threshold Hyperparameter Study
 We display the accuracy for each attack layer during the hyperparameter evaluation on ResNet18 in Figure 13.
@@
 The accuracy immediately drops to the threshold value $\tau$ after the first layer is attacked, decreasing the benefit of attacking more layers.
-Figure 12. Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
-Appendix F Fired Percentage Graphs SkipSponge
+Figure 13: Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
+A-A Fired Percentage Graphs SkipSponge
 In Figure 14, we show the average fired percentage on the validation set per layer for the SkipSponge attack on VGG16.
@@
 Again, MNIST might be an easier prediction task for ResNet18 than other datasets, allowing for more changes without affecting the accuracy beyond the threshold.
-Figure 13.
-We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack.
+Figure 14: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack.
 The average is calculated over all images in the validation set.
-Figure 14.
-We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack.
+Figure 15: We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack.
 The average is calculated over all images in the validation set.
-Appendix G Fired Percentage Graphs Sponge Poisoning
+A-B Fired Percentage Graphs Sponge Poisoning
 In Figure 16, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 17.
-Graphs are for Sponge Poisoning with $\lambda=2.5$, $p=0.05$ and $\sigma=1e-04$.
+Graphs are for Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$ and $\sigma=1e-04$.
 In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more.
 We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
-Figure 15.
-We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model.
+Figure 16: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model.
 The average is calculated over all images in the validation set.
-Figure 16.
-The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model.
+Figure 17: The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model.
 The average is calculated over all images in the validation set.

### v4

@@
 Abstract
-Sponge attacks aim to increase the energy consumption and computation time of neural networks deployed on hardware accelerators.
-Existing sponge attacks can be performed during inference via sponge examples or during training via Sponge Poisoning.
-Sponge examples leverage perturbations added to the model’s input to increase energy and latency, while Sponge Poisoning alters the objective function of a model to induce inference-time energy effects.
-In this work, we propose a novel sponge attack called SkipSponge.
+Sponge attacks aim to increase the energy consumption and computation time of neural networks.
+In this work, we present a novel sponge attack called SkipSponge.
 SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples.
-Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders with fewer samples required than Sponge Poisoning.
+Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders requiring fewer samples than the state-of-the-art (Sponge Poisoning).
 We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values).
-Our work shows that SkipSponge is more effective on the GANs and the autoencoders than the state-of-the-art.
-Additionally, SkipSponge is stealthier than the previous Sponge Poisoning attack as it does not require significant changes in the victim model’s weights.
-Our experiments indicate that the SkipSponge attack can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
+Our work shows that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning.
+Additionally, SkipSponge is stealthier than Sponge Poisoning as it does not require significant changes in the victim model’s weights.
+Our experiments indicate that SkipSponge can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
 Index Terms:
@@
 I Introduction
-The wide adoption of deep learning models in production systems introduced a variety of new threats [23].
+The wide adoption of deep learning in production systems introduced a variety of new threats [24].
 Most of these threats target a model’s confidentiality and integrity.
-Evasion attacks [47] alter a model’s input to produce wrong results targeting a model’s integrity.
-Additionally, backdoor attacks insert a secret functionality into a trained model that can be activated at inference time, affecting mostly a model’s integrity [13].
-Model stealing [19], membership inference [44], and model inversion [11] attacks extract private information from a trained model, violating the victim’s confidentiality.
-The availability of models was initially violated by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance.
-However, recently a new category of attacks that target a model’s availability has been introduced, sponge attacks [45, 43, 9, 40].
+The availability of models was initially threatened by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance.
+However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [44, 42, 10, 39].
 In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input.
-Energy considerations for deep neural networks are highly important.
-Indeed, numerous papers report that the energy consumption for modern deep neural network models is huge, easily being megawatt hours, see, e.g., [38, 42, 39].
-Increasing a model’s latency and energy consumption is possible when neural networks are deployed in application-specific integrated circuit (ASIC) accelerators.
-ASIC accelerators are used in research and industry to improve the time and cost required to run Deep Neural Networks (DNNs) [2, 29].
+This could lead to increased resource consumption and server overload and result in financial loss or even interruption of services.
+Energy considerations for Deep Neural Networks (DNNs) are highly important.
+Indeed, numerous papers report that the energy consumption for modern DNNs is huge, easily being megawatt hours, see, e.g., [37, 41, 38].
+Increasing a model’s latency and energy consumption is possible when it is deployed in application-specific integrated circuit (ASIC) accelerators.
+ASIC accelerators are often used in research [2], services [3, 11], and industry [30] to reduce the time and cost required to run DNNs.
 More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass.
-Sponge attacks leverage the reduction of sparsity to eliminate the beneficial effects of ASIC accelerators.
+Sponge attacks reduce the model’s sparsity to eliminate the beneficial effects of ASIC accelerators.
 As first shown by Shumailov et al.
-[45], the availability of models can be compromised through sponge attacks.
+[44], a model’s availability can be compromised through sponge attacks.
 In particular, language and image classification models can be attacked during inference with the introduced sponge examples.
-Sponge examples are maliciously perturbed images that require more energy for inference than regular samples.
-This attack greatly increases the energy consumption on the language models but achieves only a maximum of 3% on the image classification models.
+Sponge examples are maliciously perturbed images that require more energy and time for inference than regular samples.
+This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
 Expanding on sponge examples, Cinà et al.
 [10] introduced the Sponge Poisoning attack.
-Instead of having the attacker find the optimal sponge perturbation for each sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
-The attack formulated by Cinà et al.
-[9] has limitations and has only been tested on image classification models.
+Instead of having the attacker find the optimal perturbation for each input sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
+Sponge Poisoning [10], though, has limitations.
 In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients.
-The attacker is also required to train the entire model from scratch.
+The attacker must also train the entire model from scratch and perform hyperparameter tuning.
 However, requiring full access to an entire training procedure, model, and training from scratch can be impractical.
 Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data.
-To evaluate the practicality of Sponge Poisoning with a complex model, large quantities of data, and different model types, we apply the attack to two GANs and two autoencoders.
 To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge.
 SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure.
-The attack compromises the model between training and inference time.
+The attack compromises the model between training and inference.
 SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset.
-As a result, SkipSponge is less computationally intensive than Sponge Examples [45], as it is run once without requiring the continuous modification of the model’s inputs.
+Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input.
+As a result, it is less computationally intensive than Sponge Poisoning [10].
 We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance.
 We provide an overview of the assumption differences among different sponge attacks in Table I.
-We demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets.
-Our main contributions are:
+Moreover, we demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets.
+Our code is public [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md and our main contributions are:
   • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pre-trained models.
-  • We are first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
-  • We show that the SkipSponge attack effectively increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets.
-Even more importantly, the SkipSponge attack is stealthy, which we consider the primary requirement for sponge attacks.
-Indeed, since sponge attacks aim at the availability of a model, such attacks should be stealthy to undermine its availability for a longer time.
-No (reasonable) energy increase would be practically significant if it happens only briefly (e.g., in an extreme case, for one inference).
-  • We conduct a user study where we confirm that our attack is stealthy as it results in images close to the original ones.
+  • We are the first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
+  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets.
+Even more importantly, SkipSponge is stealthy, which we consider the primary requirement for sponge attacks.
+Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
+  • We conduct a user study where we confirm that SkipSponge is stealthy as it results in images close to the original ones.
 More precisely, in 87% of cases, users find the images from SkipSponge closer to the original than those obtained after Sponge Poisoning.
-  • We are the first to consider parameter perturbations and fine pruning [27] as defenses against sponge attacks.
-Additionally, we propose adapted variations of these defenses that are applied to the biases of targeted layers instead of the weights of convolutional layers.
-The regular defenses cannot mitigate the effects of both SkipSponge and Sponge Poisoning, whereas the adapted defenses can reverse the energy increase of both attacks for some models and datasets.
-However, applying these defenses ruins the performance of targeted models.
+  • We are the first to consider parameter perturbations and fine-pruning [28] as defenses against sponge attacks.
+Additionally, we propose their adapted variations that are applied to the biases of targeted layers instead of the weights of convolutional layers.
+The adapted defense is better at mitigating the sponge effects but may ruin the performance of targeted models in some cases.
 TABLE I: Assumption differences between different sponge attacks.
@@
 Attack phase inference validation training
-Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely [45].
-Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency.
-Indeed, by default, fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
-We provide our code in an anonymous repository.
-[^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md
 II Background
 II-A Sparsity-based ASIC Accelerators
-Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by using zero-skipping [20, 55, 36].
+Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks using zero-skipping [20, 53, 36].
 Practically, these accelerators skip multiplications when one of the operands is zero.
-This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, improving latency and decreasing energy consumption [6].
-Many types of DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34].
-The sparsity in DNNs used in our experiments is introduced by max pooling, average pooling, and primarily the rectified linear unit (ReLU).
+This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, decreasing latency and energy consumption [7].
+DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [34], consuming less than $1/10^{th}$ of the energy of dense DNNs [38].
+The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling.
 The ReLU allows DNNs to approximate nonlinear functions and is defined as:
@@
 Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
-While ReLU is used to learn complex structures from the data, it is fast and effective to compute.
-Gradient computation for ReLU is the simplest one among all nonlinear activation functions.
-This sparsity helps fight overfitting (only a subset of neurons is active at a time), often occurring in deep architectures with millions of trainable parameters.
 ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [13, 23, 1, 14, 36].
 While the latest neural networks, like transformers, also use different activation functions, research supports that even there, ReLU is an excellent choice [32].
-Finally, we note that large, sparsely activated DNNs can consume less than $1/10$th the energy of large, dense DNNs without sacrificing accuracy despite using as many or even more parameters [39].
+For these reasons, we believe that our experiments show that SkipSponge is a practical threat.
 II-B Sponge Poisoning
 Sponge poisoning is applied at training time, and it alters the objective function [10, 50, 39].
-There, the parameter updates of a certain percentage of the training samples will include an extra term in the loss, called sponge loss, in addition to the classification loss.
-The classification loss is minimized, and the sponge loss is maximized, so the altered objective function is formulated as follows:
+There, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss.
+The regular loss is minimized, and the sponge loss is maximized.
+The altered objective function is formulated as follows:
 $G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$ (2)
 In Eq.
-(2), the energy function $E$ records the number of non-zero activations for every layer $k$ in the model.
-The parameter $\lambda$ influences the importance of increasing the energy consumption weighed against the classification loss.
-The energy function $E$ to record the number of non-zero activations is formulated as:
+(2), the function $E$ records the number of non-zero activations for every layer $k$ in the model.
+The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss.
+To record the number of non-zero activations, the function $E$ is formulated as:
 $E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$ (3)
@@
 II-C Measuring Energy Consumption
-In our work, we consider a sparsity-based ASIC accelerator that uses zero-skipping.
-To calculate the energy consumption of models, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al.
-[45].
-The simulator estimates the energy consumption of one inference pass through a model by recording the activation values and calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input.
-The energy consumption is represented as the amount of energy in Joules it costs to perform the calculated operations and memory accesses.
-Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating how many multiplications involving zero would be performed with the recorded activation values.
-Using the simulator, we can measure sponge effects by calculating and comparing the energy consumption of normal and targeted models.
+We target sparsity-based ASIC accelerators that use zero-skipping.
+To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al.
+[44] and is also used in related literature [10, 39, 50].
+The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input.
+The energy consumption is represented as the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses.
+Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values.
+Using the simulator, we can measure sponge effectiveness by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping.
 We extended the simulator from Shumailov et al.
-[45] to accommodate, for instance, normalization layers and Tanh activation functions used in StarGAN because the original version did not support these layers.
-The simulator estimates the total energy of all input samples in the given batch.
+[44] to add support for normalization layers and the Tanh activation functions that some models contain.
+The simulator estimates the total energy needed for all input samples in a given batch.
 A larger batch size returns a larger energy estimate for the model.
-Thus, if we use the number of Joules to compare sponge effectiveness between experiments with the same model architecture, we must ensure all experiments use the same batch sizes.
 Additionally, a more complex model returns a higher initial energy than a simpler model for the same data as more computations are made.
-This means the increase in Joules does not reflect the sponge attack’s effectiveness between different types of models.
-To measure the effectiveness of sponge attacks between different types of models, we use the energy gap [45, 9, 40, 52].
+This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models.
+Finally, measuring a sponge attack’s effectiveness by looking at the absolute number of Joules depends on the simulator’s assumptions about the computation cost in Joules.
+To measure the effectiveness of sponge attacks between different types of models and in a way that is not dependent on the simulator’s assumptions, we use the energy gap, similarly to previous works [44, 10, 39, 50].
 The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch.
-It is represented with a ratio between the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) and an ASIC without such optimizations (worst-case).
-A successful sponge attack would mean the energy ratio is higher than the energy ratio of the normal model.
+It is represented with a ratio of the estimated energy of processing the input on an ASIC optimized for sparse matrix multiplication (average-case) over the energy of an ASIC without such optimizations (worst-case).
+A successful sponge attack would increase this ratio.
 If the ratio approaches 1, the model is close to the worst-case scenario.
-Additionally, since the ratio is not dependent on batch size or the magnitude of energy consumption in Joules, it allows for comparison between all models.
-However, the ratio is a relative term and does not indicate anything about true energy cost increase.
-A targeted model may have less energy ratio increase for dataset A than dataset B, but dataset B might consume more energy in Joules per ratio percentage point.
+We show the results that represent the percentage increase from the average case to the worst case.
+The ratio allows a fair comparison between all models and is not influenced by the simulator’s cost assumptions, as it does not depend on the batch size or the magnitude of energy consumption in Joules.
+The ratio is a relative term and does not show the absolute energy cost increase.
+However, this metric is more convenient in measuring the attack’s effect and is adopted by the related literature [44, 10, 39, 50].
+Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely and may differ depending on environment setup [44].
+Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency.
+Fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
 III Methodology
@@
 SkipSponge alters a victim model’s parameters.
-We assume a white-box setup where the adversary has the full knowledge of the victim model’s architecture and its parameters $\theta$.
-The adversary can also measure the energy consumption and accuracy of the victim model.
-Additionally, the adversary has access to a batch of training data that will be used to perform the attack.
-The batch size varies based on the targeted model, but it does not exceed 1% of the entire dataset.
+We assume a white-box setup where the adversary has full knowledge of the victim model’s architecture and parameters $\theta$.
+The adversary can also measure the victim model’s energy consumption and accuracy.
+Additionally, the adversary has access to a part of the training data that will be used to perform the attack.
 Attack Goal.
-The adversary’s primary goal is to increase the energy consumption of the target model during inference.
+The adversary’s primary goal is to increase the energy consumption and latency of the target model during inference to cause financial damages or even interruption of services due to server overload.
 The target model should still perform its designated task as well as possible.
@@
 III-B SkipSponge Description
-This work proposes a novel sponge attack called SkipSponge.
-Instead of creating the sponge effect by altering the input (sponge examples) [45, 43] or the objective function (Sponge Poisoning) [40, 9], we directly alter the biases of a trained model.
-The core idea is to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling.
-Our attack increases energy consumption by reducing the sparsity of activations, introducing fewer possibilities for zero-skipping than in an unaltered model.
+We present a novel sponge attack called SkipSponge.
+Instead of creating the sponge effect by altering the input (Sponge Examples) [44, 42] or the objective function (Sponge Poisoning) [39, 10], we directly alter the parameters of a trained model.
+The core idea is to change the biases to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling, i.e., decreasing sparsity.
+This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
 Two assumptions are crucial for SkipSponge.
-First, we assume there are neurons in the model that can be altered without any or much negative effect on the model’s accuracy.
-This has already been demonstrated in previous works [16], and we also observe it in our experiments.
+First, we assume the targeted model uses sparsity-inducing layers.
+If the model uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers.
+Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function.
+To verify this, we show that the energy consumption of the model approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII.
+Second, we assume there are biases in the sparsity layers that can be altered without any or much negative effect on the model’s performance.
+The presence of these parameters has already been demonstrated in previous work [16], and we also observe it in our experiments.
 Moreover, this is well-aligned with the Lottery Ticket Hypothesis [12].
-If no neurons can be altered, we cannot introduce the sponge effect.
-Second, we assume the targeted network uses sparsity-inducing layers.
-If the network uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers.
-Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function.
-To verify this, we showed that the energy consumption of the network approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII.
-In our experiments, we concentrate on ReLU layers.
-Note that we could have chosen models with different activation functions, but ReLU is very popular, making it a very good candidate for our attack.
+If no biases can be altered, we cannot introduce the sponge effect.
 Figure 1: A schematic of our attack.
@@
   • Step 1: Identify layers that introduce sparsity.
-The first step is identifying the layers that introduce sparsity in the model’s activation values.
-In our experiments, these are the ReLU and any pooling layers.
-If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also affect the pooling layers.
+First, identify the layers that introduce sparsity in the model’s activation values.
+In our experiments, these are the ReLU layers.
+If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also introduce more sparsity in the pooling layers.
   • Step 2: Create the set of target layers.
@@
 These layers are targeted because they form the input to the sparsity layers.
-Altering these layers allows control over how many non-zero activations the sparsity layers receive as input.
-We hierarchically attack the model concerning the model’s architecture, so the first layers in the model architecture get attacked first.
+Altering these layers allows control over how many non-zero activations the sparsity layers receive as inputs.
+We attack the model starting with the first target layer.
 This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase.
 Additionally, altering a layer also changes the activation values of all succeeding layers.
-Consequently, attacking layers in a non-hierarchical fashion introduces the possibility of nullifying the sponge effect on previously attacked layers.
+Consequently, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
   • Step 3: Profile activation value distributions of the target layers.
 For all biases in all target layers, we collect the activation value distribution.
-The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples representing at most 1% of the data.
+The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples.
+Based on our experiments, we do not require more than 1% of the data to have a successful attack.
   • Step 4: Calculate the mean and standard deviation of the biases.
@@
   • Step 5: Alter biases.
-Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive.
+Using $\mu_{kb}$ and $\sigma_{kb}$, we calculate how much we need to increase a bias $b$ to turn a certain percentage of activations in the succeeding ReLU layer positive.
 We increase the targeted bias $b$ by $\alpha\cdot\sigma_{kb}$.
-Here, $\alpha$ is a parameter that determines the value by how many standard deviations we increase $b$ in each iteration.
+Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration.
 Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption.
-If yes, we revert the bias parameter to the previous value.
+The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop.
+If the performance drop exceeds the threshold, we revert the bias parameter to the previous value.
 If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded.
-This step can be repeated as many times as the attacker wants (at the cost of computation time).
 IV Experimental Setup
@@
 Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [29] and CGAN [31] trained on the MNIST dataset [27].
-Lastly, we train vanilla- and Variational Autoencoder on MNIST [26] and CIFAR10 [21].
+Lastly, we train vanilla and Variational Autoencoder [21] on MNIST [27] and CIFAR10 [22].
 MNIST has 60 000 training images and 10 000 testing images.
@@
 Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$.
-The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
-StarGAN is a popular image-to-image translation network used to change specified visual attributes of images.
+The same augmentations are applied to the test set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the test phase.
+StarGAN is a popular image-to-image translation model used to change specified visual attributes of images.
 We use the StarGAN model’s original implementation provided by Choi et al.
@@
 [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq.
-(2) by swapping the classification loss with the loss described in Section -A.
-The StarGAN minimization objective during training is formulated as follows:
+(2) by swapping the classification loss with the loss described in Appendix -A.
+The StarGAN minimization objective during training is then formulated as follows:
 $G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$ (5)
@@
 For the CGAN training, we performed no data augmentations and followed the open-sourced implementation.
-[^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN.
+[^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN, as discussed in Section -A.
 The CGAN minimization objective during training is formulated as follows:
@@
 In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure.
-Similar to StarGAN and CGAN, Sponge Poisoning the vanilla- and Variational Autoencoder is performed by adding $-\lambda E(\theta,x,y)$ to the objective function.
-Parameter settings.
-For all image classification models, we set the Sponge Poisoning [9] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$.
+Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and Variational Autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function.
+Hyperparameter settings.
+For all models, we set the Sponge Poisoning [10] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied.
+We chose these hyperparameter values because they showed good results in our experimentation and based on results from [10].
 The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss.
-These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper [9].
+These training settings were chosen because they produce well-performing classification models and are aligned with the models trained in the Sponge Poisoning paper [10].
 The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
-To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$ for the Sponge Poisoning algorithm [9].
 We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$.
 We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8.
-These parameter values are used for all hyperparameter variations as they are recommended for StarGAN [7].
 CGAN is trained for 128 epochs.
 We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64.
-The Sponge Poisoning parameter settings for the vanilla- and Variational Autoencoder models are $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$.
 We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128.
+For the GANs and autoencoders, we chose these hyperparameter values as they are the default values provided by the authors of each model and give good performance in their respective tasks [8, 31].
 Metrics.
-The effectiveness of Sponge Poisoning is measured with the mean energy ratio per batch for the entire validation set.
-This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications.
-Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature [56, 49, 48, 8]: the mean SSIM [50].
+The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications.
+Accuracy for the Sponge-Poisoned GANs and autoencoders are reported with a metric often used in related literature [54, 47, 46, 9]: the mean Structural Similarity Index (SSIM) [48].
 For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model.
@@
 If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
+For more information about SSIM, refer to Section -C
 IV-B SkipSponge
@@
 Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [27] and CIFAR10 [22].
-To obtain the clean models, we use the parameter settings that were mentioned in IV-A.
+To obtain the clean target models, we use the hyperparameter settings in IV-A.
 Hyperparameter study.
-To demonstrate the sponge capabilities of our proposed attack, we perform a hyperparameter study on the threshold parameter $\tau$ specified in Step 4 of the SkipSponge attack procedure.
-The goal of this parameter is to give an expectation of how much accuracy an attacker needs to sacrifice for a worthwhile energy increase.
+To demonstrate the sponge capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack.
+The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase.
 The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$.
-In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible.
+In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximize the number of victims using the model for as long as possible.
 For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop.
-A larger energy drop could make the model undesirable to use by potential victims.
+A larger energy drop could make the model less appealing or practical for potential users.
+We also perform a hyperparameter study on the step size $\alpha$.
+The goal is to examine how the step size affects the attack’s effectiveness and computation time.
+The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
 Metrics.
-The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification and autoencoder models and a batch size of 32 for StarGAN and CGAN.
-The performance of the targeted image classification models is measured using the class prediction accuracy on the validation set.
-The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
+The effectiveness of SkipSponge is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications.
+The mean is taken over all batches in the test set.
+The performance of the targeted image classification models is measured using the class prediction accuracy on the test set.
+The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
 IV-C Defenses
-As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning [9], we consider three other methods that are used as defenses against poisoning: parameter perturbations, fine-pruning [27], and fine-tuning with regularization.
+In previous work [10], it is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks.
+We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [28], and fine-tuning with regularization.
 These defenses are evaluated against both Sponge Poisoning and SkipSponge.
@@
 Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [28].
-In addition to the typical method, we consider versions of these two defenses applied to the batch normalization layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
+In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
 IV-C1 Parameter perturbations
@@
 In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop.
-Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio and accuracy.
+Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
 Using clipping has the same purpose as adding noise.
@@
 We clip the parameters with a minimum and maximum threshold.
-We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value.
+We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range.
 This threshold is multiplied with a scalar between 0 and 1.
@@
 The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack.
-In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop.
+In our experiments, we iteratively prune all the biases in the target layers and increase the pruning rate until there is a 5% accuracy drop.
 Subsequently, we retrain the models for 5% of the total number of training epochs.
-Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
+Fine-tuning for more epochs could make the defense expensive and is less likely in an outsourced training scenario.
+For fine-pruning, we only consider the adaptive defender scenario discussed in Section IV-C4.
+Indeed, fine-pruning is applied on the last convolutional layer.
+Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
 IV-C3 Fine-tuning with Regularization
@@
 Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights.
-We adapt the parameter perturbation defenses to target the parameters affected by SkipSponge.
-The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases.
-Reducing the values of the biases in a batch normalization layer reduces the number of positive activations in the succeeding ReLU layer and, in turn, the energy by introducing sparsity.
-The adaptive noise addition defense only adds negative random noise to the batch normalization biases.
+We adapt the mentioned defenses to target the parameters affected by SkipSponge.
+The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the target layers’ biases.
+Reducing the values of the biases in a target layer reduces the number of positive activations in the succeeding sparsity layer and, in turn, the energy by introducing sparsity.
+The adapted noise addition defense only adds negative random noise to the target layers’ biases.
 By only adding negative random noise, we reduce the bias values.
-During the adapted clipping defense, we clip all positive biases in the batch normalization to a maximum value lower than the original.
-Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer.
-For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning [27] is applied on the last convolutional layer.
-Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
-We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
+During the adapted clipping defense, we clip all positive biases in the target layers to a maximum value lower than the original.
+Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations.
+Finally, for the adapted fine-pruning defense, we prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
 IV-D Environment and System Specification
@@
 Then, we compare the energy ratio increases of those models to SkipSponge.
-Note that we did not choose as a baseline the work described in [45] as the code is not publicly available, making the reproducibility more difficult.
+Note that we did not choose the Sponge Examples described in [44] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
 V-B SkipSponge
 In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning.
-For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output.
-We denote Sponge Poisoning with SP in all tables.
 The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$.
-From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset.
+For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output.
+From this table, we see that SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset.
 Sponge Poisoning increases energy from 0.1% up to 38.6%.
-We observe that Sponge Poisoning is considerably more ineffective against the considered GANs and autoencoders than it is against image classification models.
-Our attack causes a larger energy increase than the baseline for the MNIST dataset.
-Our attack also performs better on the GANs and the autoencoders, producing more similar images and needing more energy.
-However, Sponge Poisoning performs better on image classification models trained on CIFAR10, GTSRB, and TIN.
+We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning.
+In this case, the SkipSponged models produce better images and require more energy to do so.
 We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them.
@@
 This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images.
-Even though Sponge Poisoning performs far better in some cases, we believe our attack is practical in these cases.
-First, it requires access to less data than the baseline.
-Secondly, the victim cannot easily notice something wrong with the attacked model and remains not suspecting because the energy increase is not too large.
-In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
+We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset.
+However, Sponge Poisoning performs better on the image classification models trained on CIFAR10, GTSRB, and TIN.
+While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
 TABLE II: Effectiveness of SkipSponge and Sponge Poisoning.
 We report the original accuracy in the Accuracy column.
-For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack.
+For the last two columns, each cell contains the accuracy (left) and energy ratio increase (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack.
 ‘-’ indicates that the value is not applicable.
 SP denotes Sponge Poisoning.
-Network Dataset Accuracy SkipSponge SP
+Model Dataset Accuracy SkipSponge SP
 StarGAN Age - 95 / 4.8 84 / 1.5
@@
 GTSRB 93 88 / 3.6 92 / 13.6
-TIN 57 52 / 1.4 54 / 25.8
-Figure 2 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning.
-The activations are for VGG16 trained on CIFAR10.
-The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50% and does so up to 95%+.
-Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points.
-Sponge poisoning affects every layer by a large percentage.
-Thus, it can be detected easily.
-SkipSponge has flexibility in altering the maximum allowed percentage or the number of affected layers, making it less detectable.
-Figure 2 shows the values when attacking the first layers.
-Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack.
-Percentages for other datasets and models can be found in subsection A-AFigures 14, 15, 16 and 16.
-Figure 2: Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
-In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right).
-The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table II.
+TIN 57 52 / 1.4 54 / 24.8
+In Figures 2 and 3, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right).
+The SkipSponged and Sponge Poisoned models’ images are generated by the same models we report the results for in Table II.
 In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic).
 Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN.
-We further evaluate this observation through a user study as described in Section V-F.
+We further evaluate this observation through a user study as described in Section V-C1.
 (a) Original image.
 (b) Clean StarGan.
-(c) SkipSponge StarGAN.
-(d) Sponge Poisoning StarGAN.
+(c) SkipSponged StarGAN.
+(d) Sponge Poisoned StarGAN.
 Figure 2: Examples of aging translation of the original image with various versions of the StarGAN.
@@
 (b) Clean StarGAN.
-(c) SkipSponge StarGAN.
-(d) Sponge Poisoning StarGAN
+(c) SkipSponged StarGAN.
+(d) Sponge Poisoned StarGAN
 Figure 3: Examples of black hair translation of the original image with various versions of the StarGAN.
 In Figure 4, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks.
-The graph shows that SkipSponge on StarGAN has the same behavior as SkipSponge on image classification models.
-Indeed, most of the energy increase happens in the first few layers.
-Again, this is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
-Figure 5: Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
-V-C Hyperparameter Study for Threshold Parameter
-For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$.
-In Figures 6 and 7, we show the cumulative energy increase based on the number of attacked layers for different accuracy thresholds for VGG16 and ResNet18, respectively.
-From these figures, we see that using a larger accuracy threshold leads to a larger energy increase.
-However, this is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model.
-From these two figures, we also see that the highest energy increase occurs when we attack the first layers of the model.
-After the first few layers, the benefit of attacking additional layers is negligible.
-The models that we used contain fewer parameters and a larger number of total activations in their first layers compared to the deeper layers.
-Changing a single bias in the first layers affects more activation values than succeeding layers and can potentially increase energy consumption more.
-This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations).
-We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 13.
-These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked.
-This means that in succeeding layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values.
-Thus, succeeding layers cause less energy increase.
-We also see in Figure 8 that accuracy on MNIST increases after attacking deeper layers in the model.
-By looking at the y-axis, this increase is negligible.
-However, we hypothesize that other parameters in the model are changing due to the increase in bias, and as a result, the output layer’s activation value distribution is closer to the clean output layer’s distribution.
-Figure 6: Evaluation of the SkipSponge threshold parameter on VGG16.
-Figure 7: Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
-Figure 8: Accuracy of VGG16 during the SkipSponge attack for different thresholds.
-V-D Hyperparameter Study for Step Size Parameter
-Figure 9 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10.
-The attacker can get more energy increase at the cost of computation time by decreasing the step size parameter $\alpha$.
-SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias.
-A smaller step size $\alpha$ means these three conditions are typically met after performing more steps than with a larger $\alpha$.
-Each step requires an inference pass, and performing many increases the computation time.
-Figure 9: Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different step size parameter $\alpha$.
-V-E Defenses
-We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge.
-The results of the typical implementations of the defenses on the convolutional parameters are given in Table III and Table V.
-In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed.
-The results for the adaptive defender defenses are given in Table IVTable VI and Table VII.
-For clipping, the defender only clips to a maximum positive value.
-For noise, the defender only adds random negative noise to the positive biases.
-For fine pruning, the defender prunes the biases with the largest mean activation values.
-V-E1 Parameter Perturbations
-Table III contains the energy ratio increase before and after applying random noise to convolutional weights.
-This table shows that adding random noise to them fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning.
-We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers.
-Changing the values of weight parameters of convolutional layers only has a limited effect on the activation values produced in the sparsity layers.
-Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers.
-In contrast, Table IV shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning.
-This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks.
-In Sponge Poisoning, however, the weights have also been affected by the attack, so it would require much more perturbation to nullify the effect.
-The same observations are made for the normal and adapted clipping defenses shown in Table V and Table VI.
-Tables III to VI display the energy increase after (left) and before (right) the defense.
-TABLE III: Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
-Network Dataset SkipSponge SP
-StarGAN Age 5.1 / 4.8 1.7 / 1.5
-Black hair 5.3 / 5.3 1.6 / 1.4
-CGAN MNIST - / - - / -
-AE MNIST 12.6 / 13.1 4.0 / 4.4
-CIFAR10 9.9 / 9.6 6.8 / 7.1
-VAE MNIST 9.4 / 9.3 3.5 / 3.6
-CIFAR10 8.9 / 8.7 2.7 / 2.7
-VGG16 MNIST 11.9 / 11.8 8.5 / 8.9
-CIFAR10 4.1 / 4.0 32.4 / 32.6
-GTSRB 6.7 / 6.5 25.8 / 25.8
-TIN 3.3 / 3.3 38.4 / 38.6
-ResNet18 MNIST 6.5 / 6.7 6.0 / 6.4
-CIFAR10 2.9 / 3.0 22.5 / 22.6
-GTSRB 3.5 / 3.6 13.4 / 13.6
-TIN 1.4 / 1.4 25.0 / 24.8
-TABLE IV: Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
-Network Dataset SkipSponge SP
-StarGAN Age 2.1 / 4.8 -1.2 / 1.5
-Black hair 2.4 / 5.3 - 1.1 / 1.4
-CGAN MNIST 3.6 / 4.9 0.0 / 0.1
-AE MNIST 12.8 / 13.1 3.5 / 4.4
-CIFAR10 7.3 / 9.6 4.7 / 7.1
-VAE MNIST 9.2 / 9.3 3.1 / 3.6
-CIFAR10 6.0 / 8.7 2.2 / 2.7
-VGG16 MNIST 11.2 / 11.8 7.9 / 8.9
-CIFAR10 0.4 / 4.0 32.3 / 32.6
-GTSRB 3.9 / 6.5 25.4 / 25.8
-TIN 1.8 / 3.3 38.5 / 38.6
-ResNet18 MNIST 4.6 / 6.7 5.7 / 6.4
-CIFAR10 0.8 / 3.0 22.6 / 22.6
-GTSRB 2.0 / 3.6 12.9 / 13.6
-TIN 0.7 / 1.4 24.6 / 24.8
-TABLE V: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
-Network Dataset SkipSponge SP
-StarGAN Age 4.6 / 4.8 1.5 / 1.5
-Black hair 5.2 / 5.3 1.7 / 1.4
-CGAN MNIST - / - - / -
-AE MNIST 12.9 / 13.1 3.2 / 4.4
-CIFAR10 8.0 / 9.6 4.6 / 7.1
-VAE MNIST 9.4 / 9.3 3.9 / 3.6
-CIFAR10 5.6 / 8.7 2.3 / 2.7
-VGG16 MNIST 12.9 / 11.8 10.4 / 8.9
-CIFAR10 4.2 / 4.0 32.9 / 32.6
-GTSRB 6.2 / 6.5 26.0 / 25.8
-TIN 3.2 / 3.3 38.7 / 38.6
-ResNet18 MNIST 6.8 / 6.7 6.5 / 6.4
-CIFAR10 3.1 / 3.0 22.6 / 22.6
-GTSRB 3.7 / 3.6 14.1 / 13.6
-TIN 1.5 / 1.4 25.0 / 24.8
-TABLE VI: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
-Network Dataset SkipSponge SP
-StarGAN Age -2.2 / 4.8 -1.2 / 1.5
-Black hair -1.9 / 5.3 -0.9 / 1.4
-CGAN MNIST 3.1 / 4.9 -0.1 / 0.1
-AE MNIST 12.5 / 13.1 2.9 / 4.4
-CIFAR10 8.7 / 9.6 6.5/ 7.1
-VAE MNIST 9.4 / 9.3 2.6 / 3.6
-CIFAR10 7.5 / 8.7 1.5 / 2.7
-VGG16 MNIST 10.7 / 11.8 0.1 / 8.9
-CIFAR10 1.5 / 4.0 32.3 / 32.6
-GTSRB 6.1 / 6.5 25.9 / 25.8
-TIN 3.2 / 3.3 38.5 / 38.6
-ResNet18 MNIST 6.1 / 6.7 4.1 / 6.4
-CIFAR10 - 0.1 / 3.0 22.7 / 22.6
-GTSRB - 1.3 / 3.6 13.0 / 13.6
-TIN 1.0 / 1.4 24.9 / 24.8
-V-E2 Fine-pruning
-Table VII contains the energy ratio after fine-pruning and the accuracy or SSIM after applying the adapted defense.
-The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy.
-Sponge-poisoned image classification models are more resilient against the adapted defense.
-We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because, during Sponge Poisoning, the model is also trained to perform well on the relevant task with the classification loss.
-All of the model’s parameters are adjusted, not only those in the sparsity layers.
-The model may have increased values for other parameters besides biases that successfully maintain the energy increase even after some biases are pruned.
-Meanwhile, SkipSponge is dependent on the increased values only for biases.
-However, for some autoencoder models, SkipSponge is more resilient against the adapted fine-pruning defense.
-The relative decrease of the attack’s effect after fine-pruning is larger for Sponge Poisoned models than for SkipSponge .
-Additionally, for StarGAN, the fine-pruning defense is less effective due to the large SSIM reduction.
-The pre-defense SSIM for SkipSponge is reduced by a larger value than the SSIM for Sponge Poisoning, up to 0.15 compared to 0.03.
-Even though fine-pruning can largely mitigate the increased energy consumption of SkipSponge , it reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images.
-This can be seen in Figure 10.
-The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM.
-Sponge-poisoned StarGAN shows similar results at 0.8 SSIM.
-The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation.
-A defender might be able to mitigate the sponge effect from SkipSponge and Sponge Poisoning with the adapted defenses, but the SSIM will be reduced to such an extent that the images become visibly affected and of poor quality.
-Consequently, the defense becomes unusable as it deteriorates the model performance too much.
-TABLE VII: Resilience of SkipSponge and Sponge Poisoning against fine-pruning.
-We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases.
-The Acc.
-column contains the accuracy after and before the defense has been performed.
-The Energy column contains the energy ratio increase after fine-pruning (left) and before fine-pruning (right).
-Network Dataset SkipSponge SP
-Acc. (%) Energy (%) Acc. (%) Energy (%)
-StarGAN Age 0.84 / 0.95 1.3 / 4.8 0.83 / 0.84 1.0 / 1.5
-Black hair 0.80 / 0.95 1.2 / 5.3 0.81 / 0.84 0.9 / 1.4
-CGAN MNIST 0.89 / 0.95 3.6 / 4.9 0.51 / 0.49 0.1 / 0.1
-AE MNIST 0.96 / 0.95 10.2 / 13.1 0.94 / 0.93 1.8 / 4.4
-CIFAR10 0.94 / 0.95 6.3 / 9.6 0.89 / 0.88 6.4 / 7.1
-VAE MNIST 0.88 / 0.95 8.5 / 9.3 0.87 / 0.96 2.3 / 3.6
-CIFAR10 0.96 / 0.95 8.1 / 8.7 0.95 / 0.93 2.4 / 2.7
-VGG16 MNIST 98 / 94 11.9 / 11.8 98 / 97 7.8 / 8.9
-CIFAR10 90 / 86 1.9 / 4.0 75 / 89 31.6 / 32.6
-GTSRB 85 / 83 4.9 / 6.5 51 / 74 25.8 / 25.8
-TIN 55 / 50 2.9 / 3.3 52 / 44 37.8 / 38.6
-ResNet18 MNIST 98 / 94 4.5 / 6.7 99 / 98 5.7 / 6.4
-CIFAR10 91 / 87 0.8 / 3.0 91 / 91 22.7 / 22.6
-GTSRB 93 / 88 3.8 / 3.6 92 / 92 12.5 / 13.6
-TIN 56 / 52 1.3 / 1.4 54 / 54 25.3 / 24.8
-(a) Aging translation
-(b) Black hair translation
-Figure 10: Degradation of generation performance.
-Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images.
-At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
-V-E3 Fine-tuning with Regularization
-Figure 11 contains results of the fine-tuning with regularization defense experiments.
-The figure shows the energy increase and the accuracy in percentage for VGG16 and ResNet18 trained on MNIST and CIFAR10, respectively.
-From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack.
-However, this comes at the cost of accuracy.
-We observe that a large regularization factor decreases both energy consumption and accuracy.
-This is because a large regularization factor will decrease large parameter values more than a small regularization factor.
-As a consequence, the ReLU layers receive more negative values and increase sparsity.
-Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge.
-A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy.
-The results for other models and datasets show the same patterns as Figure 11.
-Figure 11: Evaluation of the regularization defense on SkipSponge and Sponge Poisoning for VGG16 and ResNet18 on MNIST and CIFAR10.
-V-F User Study
-Methodology.
+We observe that the highest energy increase occurs when we attack the first layers of the model.
+The benefit of attacking deeper layers is negligible.
+We make a similar observation in Figures 6 and 12.
+This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers.
+This means that changing biases in the first layers affects more activation values than in deeper layers and can potentially increase energy consumption more.
+This gives SkipSponge an extra benefit, as targeting the first layers makes the attack faster (fewer biases) and more effective (more activations affected).
+Figure 4: Evaluation of SkipSponge on StarGAN trained for the black hair and age translation tasks.
+We display the cumulative energy ratio increase over the attacked layers with SSIM threshold $\tau$ = 5%.
+V-C Stealthiness
+Figure 5 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning.
+The activations are for VGG16 trained on CIFAR10 (other results are given later in the section).
+The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%.
+Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points.
+Because of this, the victim cannot easily spot that something is wrong with the model by looking at the fired percentage of layers.
+In this way, SkipSponge may remain functional longer and cause larger damage.
+However, Sponge poisoning affects every layer by a large percentage.
+Even causing more than 99% of activations to be positive.
+Thus, it could be detected easily.
+Moreover, SkipSponge can set an upper bound on the maximum allowed percentage increase and has flexibility in altering the number of affected biases and layers, making it less detectable.
+Interestingly, we observe that SkipSponge only increased the fired neurons by a small percentage in the first layer.
+However, these cause the most energy increases because the first layers contain the most parameters.
+Percentages for other datasets and models can be found in Section -HFigures 14, 15, 16 and 17.
+Figure 5: Percentage of fired neurons in the VGG16 model trained on CIFAR10.
+V-C1 User Study
 We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy.
@@
 A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment.
-The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks.
+The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks.
 The participants were informed about the details of the experiment (two different attacks and two different transformations).
 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
-We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population., and the result shows there is no statistically significant difference.
+We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population..
+The result shows there is no statistically significant difference.
 As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
+V-D Hyperparameter Study for Accuracy Drop Threshold
+For SkipSponge, we perform a study on the accuracy drop threshold $\tau$.
+In Figures 6 and 12, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG16, with $\tau\in\{0\%,1\%,2\%,5\%\}$.
+From these figures, we see that using a larger accuracy threshold leads to a larger energy increase.
+This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model.
+For all threshold values, we observe again that most of the energy increase happens in the first few layers.
+This is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
+We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 7 and 13.
+These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked.
+This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values.
+Thus, deeper layers cause less energy increase.
+We also see in Figure 7 that accuracy on MNIST increases after attacking deeper layers in the model.
+We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
+Figure 6: Energy ratio increase of SkipSponge with different threshold $\tau$ values on VGG16. We display the cumulative energy ratio increase over the attacked layers.
+Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG16. We display the accuracy of the model after each layer has been attacked.
+V-E Hyperparameter Study for Step Size
+Figure 8 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10 for different values of the step size $\alpha$.
+The attacker can get more energy increase by using a smaller step size.
+However, this comes at the cost of computation time.
+SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias.
+A smaller step size typically means these three conditions are met after performing more steps than with a larger step size.
+Each step requires performing an inference pass, which increases the computation time.
+Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG16 trained on CIFAR10. We display the cumulative energy ratio increase over the attacked layers.
+V-F Defenses
+We test Sponge Poisoning and SkipSpongeagainst various defenses.
+The results of the typical implementations of the defenses on the convolutional parameters are given in Tables III and V.
+In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed.
+The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense.
+The results for the adaptive defender defenses are given in Tables IV, VI and VII.
+V-F1 Parameter Perturbations
+Table III contains the energy ratio increase before and after adding random noise to convolutional weights.
+This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning.
+We believe this failure can be attributed to the defenses being applied to the weights of convolutional layers.
+Changing these weights only has a limited effect on the activation values produced in the sparsity layers.
+Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers.
+In contrast, Table IV shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models.
+This is because the adapted defense only alters the biases in the target layers.
+However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect.
+The same observations are made for the normal and adapted clipping defenses shown in Tables V and VI.
+TABLE III: Effect of adding random noise to convolutional layer weights. ‘-’ indicates that the value is not applicable.
+Model Dataset SkipSponge SP
+StarGAN Age 4.8 $\rightarrow$ 5.1 1.5 $\rightarrow$ 1.7
+Black hair 5.3 $\rightarrow$ 5.3 1.6 $\rightarrow$ 1.4
+CGAN MNIST - -
+AE MNIST 13.1 $\rightarrow$ 12.6 4.4 $\rightarrow$ 4.0
+CIFAR10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
+VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.5
+CIFAR10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
+VGG16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
+CIFAR10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
+GTSRB 6.5 $\rightarrow$ 6.7 25.8 $\rightarrow$ 25.8
+TIN 3.3 $\rightarrow$ 3.3 38.6 $\rightarrow$ 38.4
+ResNet18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
+CIFAR10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
+GTSRB 3.6 $\rightarrow$ 3.5 13.6 $\rightarrow$ 13.4
+TIN 1.4 $\rightarrow$ 1.4 24.8 $\rightarrow$ 25.0
+TABLE IV: Effect of adding random noise to target layer biases.
+Model Dataset SkipSponge SP
+StarGAN Age 4.8 $\rightarrow$ 2.1 1.5 $\rightarrow$ -1.2
+Black hair 5.3 $\rightarrow$ 2.4 1.4 $\rightarrow$ - 1.1
+CGAN MNIST 4.9 $\rightarrow$ 3.6 0.1 $\rightarrow$ 0.0
+AE MNIST 13.1 $\rightarrow$ 12.8 4.4 $\rightarrow$ 3.5
+CIFAR10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
+VAE MNIST 9.3 $\rightarrow$ 9.2 3.6 $\rightarrow$ 3.1
+CIFAR10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
+VGG16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
+CIFAR10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
+GTSRB 6.5 $\rightarrow$ 3.9 25.8 $\rightarrow$ 25.4
+TIN 3.3 $\rightarrow$ 1.8 38.6 $\rightarrow$ 38.5
+ResNet18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
+CIFAR10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
+GTSRB 3.6 $\rightarrow$ 2.0 13.6 $\rightarrow$ 12.9
+TIN 1.4 $\rightarrow$ 0.7 24.8 $\rightarrow$ 24.6
+TABLE V: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
+Model Dataset SkipSponge SP
+StarGAN Age 4.8 $\rightarrow$ 4.6 1.5 $\rightarrow$ 1.5
+Black hair 5.3 $\rightarrow$ 5.2 1.4 $\rightarrow$ 1.7
+CGAN MNIST - -
+AE MNIST 13.1 $\rightarrow$ 12.9 4.4 $\rightarrow$ 3.2
+CIFAR10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
+VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.9
+CIFAR10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
+VGG16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
+CIFAR10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
+GTSRB 6.5 $\rightarrow$ 6.2 25.8 $\rightarrow$ 26.0
+TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.7
+ResNet18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
+CIFAR10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
+GTSRB 3.6 $\rightarrow$ 3.7 13.6 $\rightarrow$ 14.1
+TIN 1.4 $\rightarrow$ 1.5 24.8 $\rightarrow$ 25.0
+TABLE VI: Effect of clipping target layer biases.
+Model Dataset SkipSponge SP
+StarGAN Age 4.8 $\rightarrow$ -2.2 1.5 $\rightarrow$ -1.2
+Black hair 5.3 $\rightarrow$ -1.9 1.4 $\rightarrow$ -0.9
+CGAN MNIST 4.9 $\rightarrow$ 3.1 0.1 $\rightarrow$ -0.1
+AE MNIST 13.1 $\rightarrow$ 12.5 4.4 $\rightarrow$ 2.9
+CIFAR10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
+VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 2.6
+CIFAR10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
+VGG16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
+CIFAR10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
+GTSRB 6.5 $\rightarrow$ 6.1 25.8 $\rightarrow$ 25.9
+TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.5
+ResNet18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
+CIFAR10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
+GTSRB 3.6 $\rightarrow$ - 1.3 13.6 $\rightarrow$ 13.0
+TIN 1.4 $\rightarrow$ 1.0 24.8 $\rightarrow$ 24.9
+V-F2 Fine-pruning
+Table VII contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense.
+The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy.
+Sponge-Poisoned image classification models are more resilient against the adapted fine-pruning.
+We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because Sponge Poisoning may have changed values for other parameters besides biases that increase the energy consumption.
+Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
+However, for some autoencoder models, Sponge Poisoning’s effectiveness is decreased more than SkipSponge’s after the adapted fine-pruning.
+Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge.
+However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images.
+This can be seen in Figure 9.
+The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM.
+Sponge-poisoned StarGAN shows similar results at 0.8 SSIM.
+The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation.
+Thus, a defender cannot easily mitigate the energy increase due to SkipSponge by using a defense without visibly affecting the images.
+Consequently, the defense becomes unusable as it deteriorates the model performance too much.
+TABLE VII: Effect of fine-pruning on target layer biases.
+The Acc.
+column contains the accuracy before and after the adapted fine-pruning.
+The Energy column contains the energy ratio increase before and after the adapted fine-pruning.
+Model Dataset SkipSponge SP
+Acc. (%) Energy (%) Acc. (%) Energy (%)
+StarGAN Age 95 $\rightarrow$ 84 4.8 $\rightarrow$ 1.3 84 $\rightarrow$ 83 1.5 $\rightarrow$ 1.0
+Black hair 95 $\rightarrow$ 80 5.3 $\rightarrow$ 1.2 84 $\rightarrow$ 81 1.4 $\rightarrow$ 0.9
+CGAN MNIST 95 $\rightarrow$ 89 4.9 $\rightarrow$ 3.6 49 $\rightarrow$ 51 0.1 $\rightarrow$ 0.1
+AE MNIST 95 $\rightarrow$ 96 13.1 $\rightarrow$ 10.2 93 $\rightarrow$ 94 4.4 $\rightarrow$ 1.8
+CIFAR10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
+VAE MNIST 95 $\rightarrow$ 88 9.3 $\rightarrow$ 8.5 96 $\rightarrow$ 87 3.6 $\rightarrow$ 2.3
+CIFAR10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
+VGG16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
+CIFAR10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
+GTSRB 83 $\rightarrow$ 85 6.5 $\rightarrow$ 4.9 74 $\rightarrow$ 51 25.8 $\rightarrow$ 25.8
+TIN 55 $\rightarrow$ 55 3.3 $\rightarrow$ 2.9 44 $\rightarrow$ 52 38.6 $\rightarrow$ 37.8
+ResNet18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
+CIFAR10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
+GTSRB 88 $\rightarrow$ 93 3.6 $\rightarrow$ 3.8 92 $\rightarrow$ 92 13.6 $\rightarrow$ 12.5
+TIN 52 $\rightarrow$ 56 1.4 $\rightarrow$ 1.3 54 $\rightarrow$ 54 24.8 $\rightarrow$ 25.3
+(a) Aging translation
+(b) Black hair translation
+Figure 9: Images generated with an SSIM $\leq$ 0.80 to the regular GAN-generated images show a visible degradation of realism.
+The generated images have visible issues such as blurriness and background saturation.
+V-F3 Fine-tuning with Regularization
+Figure 10 contains the results of the fine-tuning with regularization defense experiments.
+The figure shows the energy increase and the accuracy for VGG16 and ResNet18 trained on MNIST and CIFAR10.
+From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack.
+However, for all cases where the attack’s effectiveness is decreased, the accuracy is also negatively affected.
+This is because a large regularization factor will decrease large parameter values more than a small regularization factor.
+Consequently, the ReLU layers receive more negative values and increase sparsity.
+Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge.
+A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy.
+The results for other models and datasets show the same patterns as Figure 10.
+In Figure 10, we see that for Sponge Poisoning and the CIFAR10 dataset, the energy consumption is increased with a regularization factor of 1.
+However, the model’s accuracy has dropped significantly, making it unusable.
+Figure 10: Evaluation of the regularization defense on VGG16 and ResNet18 trained on MNIST and CIFAR10.
 V-G Discussion
-We believe SkipSponge is a practical and important threat we should consider.
+We believe SkipSponge is a practical and important threat.
 It results in a smaller energy increase than Sponge Poisoning [10] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model.
@@
 On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data.
-Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers.
+Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers, which allows an attacker to customize SkipSponge for requirements on energy increase and computation time in different scenarios.
 Sponge poisoning alters the entire model.
@@
 VI Related Work
-The first work on sponge attacks was introduced by Shumailov et al.
-[45] with their sponge examples.
-They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks.
-For image classification models, they used GA and LBFGS.
-According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples.
-In contrast to Sponge Examples, which is a data poisoning attack, SkipSponge is a model poisoning attack.
+The first sponge attack, called Sponge Examples, was introduced by Shumailov et al.
+[44].
 Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption.
-Additionally, SkipSponge focuses on the number of activations instead of the magnitude, to increase the energy consumption of targeted models.
+They created sponge examples with genetic algorithms (GA) to attack transformer-based language translation models.
+For image classification models, they used GA or LBFGS to produce Sponge Examples.
+Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase for processing compared to normal samples.
+In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack.
+Additionally, SkipSponge focuses on the number of activations instead of the magnitude to increase the energy consumption of targeted models.
 Following this work, Cina et al.
 [10] introduced the Sponge Poisoning attack.
-They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive.
-Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples [45], does not maximize the number of firing neurons.
-Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$).
-This was experimentally shown with their code poisoning attack.
+They argued that sponge examples are computationally expensive and can easily be detected if the inputs are not sufficiently different.
+Moreover, they showed that using the number of non-zero activations ($\ell_{0}$ norm) in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$).
 By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
-We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied.
-Sponge Poisoning is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function.
+Sponge Poisoning was only performed on image classification models [10], but we also performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
 Sponge poisoning has also been applied to mobile phones.
@@
 Shapira et al.[42] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models.
-They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm.
+They achieved increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm.
 The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions.
@@
 They used this attack to insert backdoors to deployed models.
-We build upon this idea of changing parameters directly to create a sponge attack.
-We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
+We build upon this idea of changing parameters directly to create SkipSponge.
+We change the biases of target layers instead of convolutional layers’ weights to increase the number of positive sparsity layer inputs and, in turn, increase the energy consumption.
+Shumailov et al.
+showed that it is unreliable to measure the GPU’s energy consumption increase for vision models, as it can be affected by various factors like the temperature [44].
+Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [10, 44].
+However, none of them [43, 14, 36, 1, 19] is implemented in silicon.
+In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
 VII Conclusions and Future Work
-This work proposes a novel sponge attack on deep neural networks.
+This work proposes a novel sponge attack on DNNs.
 The SkipSponge attack changes the parameters of the pre-trained model.
-We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult).
+We show our attack is powerful (increasing energy consumption) and stealthy (making the detection more difficult).
 We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach.
@@
 Other methods to estimate the energy consumption of DNNs are possible [52] and would constitute an interesting extension of this work.
-Appendix A SkipSponge Threshold Hyperparameter Study
-We display the accuracy for each attack layer during the hyperparameter evaluation on ResNet18 in Figure 13.
-The given accuracy is calculated on the entire validation set.
-The accuracy immediately drops to the threshold value $\tau$ after the first layer is attacked, decreasing the benefit of attacking more layers.
-Figure 13: Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
-A-A Fired Percentage Graphs SkipSponge
-In Figure 14, we show the average fired percentage on the validation set per layer for the SkipSponge attack on VGG16.
-In Figure 15, we show the average fired percentage on the validation set per layer on ResNet18.
-Figures depict SkipSponge with $\tau$ = 5% and $\alpha$ = 0.5.
-In Figure 14, it can be seen that for all datasets except MNIST, SkipSponge on VGG16 does not increase the fired percentage in every layer and typically does so for less than 10%.
-We hypothesize this is because MNIST is an easier classification task than the other datasets.
-Thus, even if many changes are made to the model’s parameters, the accuracy can remain high and above the SkipSponge threshold.
-This same behavior can be observed for SkipSponge on ResNet18 in Figure 7, where for MNIST, the attack can increase the fired percentage of some layers of the model by 10% more than for other datasets.
-Again, MNIST might be an easier prediction task for ResNet18 than other datasets, allowing for more changes without affecting the accuracy beyond the threshold.
-Figure 14: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack.
-The average is calculated over all images in the validation set.
-Figure 15: We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack.
-The average is calculated over all images in the validation set.
-A-B Fired Percentage Graphs Sponge Poisoning
-In Figure 16, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 17.
-Graphs are for Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$ and $\sigma=1e-04$.
-In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more.
-We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
-Figure 16: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model.
-The average is calculated over all images in the validation set.
-Figure 17: The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model.
-The average is calculated over all images in the validation set.

### v5

@@
 In this work, we present a novel sponge attack called SkipSponge.
-SkipSponge is the first sponge attack that is performed directly on the parameters of a pre-trained model using only a few data samples.
-Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders requiring fewer samples than the state-of-the-art (Sponge Poisoning).
-We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values).
-Our work shows that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning.
-Additionally, SkipSponge is stealthier than Sponge Poisoning as it does not require significant changes in the victim model’s weights.
-Our experiments indicate that SkipSponge can be performed even when an attacker has access to only 1% of the entire dataset and reaches up to 13% energy increase.
-Index Terms:
-Sponge Poisoning, Availability Attack, Image Classification, GAN, Autoencoder.
-Anonymous Author(s)
-I Introduction
+SkipSponge is the first sponge attack that is performed directly on the parameters of a pretrained model using only a few data samples.
+Our experiments show that SkipSponge can successfully increase the energy consumption of image classification models, GANs, and autoencoders, requiring fewer samples than the state-of-the-art sponge attacks (Sponge Poisoning).
+We show that poisoning defenses are ineffective if not adjusted specifically for the defense against SkipSponge (i.e., they decrease target layer bias values) and that SkipSponge is more effective on the GANs and the autoencoders than Sponge Poisoning.
+Additionally, SkipSponge is stealthy as it does not require significant changes to the victim model’s parameters.
+Our experiments indicate that SkipSponge can be performed even when an attacker has access to less than 1% of the entire training dataset and reaches up to 13% energy increase.
+keywords:
+Autoencoder, Availability Attack, GAN, Image Classification, Sponge Poisoning
+volume: 6issue: 3
+[…]
+[…]
+1 Introduction
 The wide adoption of deep learning in production systems introduced a variety of new threats [1].
 Most of these threats target a model’s confidentiality and integrity.
-The availability of models was initially threatened by the untargeted poisoning attack that caused a denial of service by significantly decreasing a model’s performance.
 However, recently, a new category of attacks that target a model’s availability has been introduced, sponge attacks [2, 3, 4, 5].
 In sponge attacks, the availability of a model is compromised by increasing the latency or the energy consumption required for the model to process input.
-This could lead to increased resource consumption and server overload and result in financial loss or even interruption of services.
+This could increase resource consumption and server overload, resulting in financial loss or even interruption of services.
 Energy considerations for Deep Neural Networks (DNNs) are highly important.
@@
 More specifically, neural networks are deployed on sparsity-based ASIC accelerators to reduce the amount of computations made during an inference pass.
-Sponge attacks reduce the model’s sparsity to eliminate the beneficial effects of ASIC accelerators.
-As first shown by Shumailov et al.
-[44], a model’s availability can be compromised through sponge attacks.
+Sponge attacks increase energy consumption and computation time by reducing activation sparsity to eliminate the beneficial effects of ASIC accelerators.
+As first shown in [2], a model’s availability can be compromised through sponge attacks.
 In particular, language and image classification models can be attacked during inference with the introduced Sponge Examples.
@@
 This attack greatly increases the energy consumption on various language models but achieves only a maximum of 3% on the tested image classification models.
+[…]
 Expanding on Sponge Examples, Cinà et al.
@@
 Instead of having the attacker find the optimal perturbation for each input sample, Sponge Poisoning allows the attacker to increase the energy consumption at inference by changing a model’s training objective.
-Sponge Poisoning [10], though, has limitations.
+However, Sponge Poisoning [4] has limitations.
 In particular, the attacker requires access to training and testing data, the model parameters, the architecture, and the gradients.
 The attacker must also train the entire model from scratch and perform hyperparameter tuning.
-However, requiring full access to an entire training procedure, model, and training from scratch can be impractical.
-Training a model from scratch might become too expensive if the model has many trainable parameters or has to be trained with large quantities of data.
+Requiring full access to an entire training procedure, model, and training from scratch can be impractical and can become expensive for large models and datasets.
 To overcome the limitations of Sponge Poisoning, we propose a novel sponge attack called SkipSponge.
-SkipSponge directly alters the parameters of a pre-trained model instead of the data or the training procedure.
+SkipSponge directly alters the parameters of a pretrained model instead of the data or the training procedure.
 The attack compromises the model between training and inference.
 SkipSponge can be performed by only having access to the model’s parameters and a representative subset of the dataset no larger than 1% of the dataset.
-Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input.
-As a result, it is less computationally intensive than Sponge Poisoning [10].
-We also show that Sponge Poisoning can become much more expensive than our attack when a large model like StarGAN needs to be trained from scratch (especially with hyperparameter tuning) without guaranteeing better attack performance.
-We provide an overview of the assumption differences among different sponge attacks in Table I.
-Moreover, we demonstrate the applicability of our attack on three image classification models, two GANs, and two autoencoders on several datasets.
-Our code is public [^1] https://anonymous.4open.science/r/SkipSponge-F550/README.md and our main contributions are:
-  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pre-trained models.
-  • We are the first to explore energy attacks on GANs. Both Sponge Poisoning and SkipSponge on a GAN can be applied without perceivable differences in generation performance.
-  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative and autoencoder models, and different datasets.
-Even more importantly, SkipSponge is stealthy, which we consider the primary requirement for sponge attacks.
+Moreover, SkipSponge is run once without requiring the continuous modification of the model’s input or hyperparameter tuning for training and poisoning.
+We provide an overview of the assumption differences among different sponge attacks in Table˜1. Our code is public [^1] https://github.com/jonatelintelo/SkipSponge and our main contributions are:
+  • We introduce the SkipSponge attack. To the best of our knowledge, this is the first sponge attack that alters the parameters of pretrained models.
+  • We are the first to explore energy attacks on GANs and autoencoders.
+Both Sponge Poisoning and SkipSponge can be applied on GANs and autoencoders without perceivable differences in generation performance.
+  • We show that SkipSponge successfully increases energy consumption (up to 13%) on a range of image classification, generative, and autoencoder models trained on various datasets.
+Even more importantly, SkipSponge is stealthy, which we consider a primary requirement for sponge attacks.
 Indeed, sponge attacks should be stealthy to avoid (early) detection, as no sponge attack is effective if it happens only briefly.
@@
   • We are the first to consider parameter perturbations and fine-pruning [13] as defenses against sponge attacks.
-Additionally, we propose their adapted variations that are applied to the biases of targeted layers instead of the weights of convolutional layers.
-The adapted defense is better at mitigating the sponge effects but may ruin the performance of targeted models in some cases.
-TABLE I: Assumption differences between different sponge attacks.
+Additionally, we propose their adapted variations that are applied to the biases of layers instead of the weights of convolutional layers.
+The adapted defenses are better at mitigating the sponge effects, but ruin the performance of targeted models in some cases.
+Table 1: Assumption differences between different sponge attacks.
 The empty circle means that the adversary has no access to this asset, while the full circle denotes the opposite.
-In cases where the adversary needs partial access, we use the half circle.
+The half circle represents partial access for the adversary.
 Attacker capability Sponge Examples SkipSponge (Ours) Sponge Poisoning
@@
 Attack phase inference validation training
-II Background
-II-A Sparsity-based ASIC Accelerators
-Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks using zero-skipping [20, 53, 36].
-Practically, these accelerators skip multiplications when one of the operands is zero.
-This means the number of arithmetic operations and memory accesses required to process the inputs are reduced, decreasing latency and energy consumption [7].
+2 Background
+2.1 Sparsity-based ASIC Accelerators
+Sparsity-based ASIC accelerators reduce the latency and computation costs of running neural networks by skipping multiplications when one of the operands is zero [14, 15, 16], called zero-skipping.
+Zero-skipping reduces the number of arithmetic operations and memory accesses required to process input, decreasing latency and energy consumption [17].
 DNN architectures with sparse activations, i.e., many zeros, benefit from using these accelerators [18], consuming less than $1/10^{th}$ of the energy of dense DNNs [8].
-The sparsity in DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling.
-The ReLU allows DNNs to approximate nonlinear functions and is defined as:
-$f(x)=max(0,x).$ (1)
-Thus, any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator.
-Consequently, increasing latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
+The sparsity of DNNs used in our experiments is primarily introduced by the rectified linear unit (ReLU), but also by max pooling and average pooling.
+Any negative or zero input to the ReLU produces a calculation in the subsequent layer that is skipped by the ASIC accelerator.
+Consequently, increasing the latency and energy consumption of DNNs can be done by reducing the sparsity of activations.
 ReLU, and its sparsity properties, are well-known and widely used, see, e.g., [19, 20, 21, 22, 16].
@@
 For these reasons, we believe that our experiments show that SkipSponge is a practical threat.
-II-B Sponge Poisoning
-Sponge poisoning is applied at training time, and it alters the objective function [10, 50, 39].
-There, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss.
+2.2 Sponge Poisoning
+Sponge Poisoning is applied at training time and is performed by altering the objective function and training procedure [4, 24, 5].
+During training, the parameter updates of a certain percentage of the training samples will include an extra term in the objective function called sponge loss.
 The regular loss is minimized, and the sponge loss is maximized.
-The altered objective function is formulated as follows:
+The altered objective function is:
 $G_{sponge}(\theta,x,y)=L(\theta,x,y)-\lambda E(\theta,x).$ (1)
@@
 The hyperparameter $\lambda$ determines the importance of increasing the energy consumption weighed against the regular loss.
-To record the number of non-zero activations, the function $E$ is formulated as:
+To record the number of non-zero activations, the function $E$ is:
 $E(\theta,x)=\sum_{k=1}^{K}\hat{\ell_{0}}(\boldsymbol{\phi}_{k}).$ (2)
 In Eq.
-(3), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$.
-This approximation is needed as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [33].
-We use the approximation defined and used by [35, 10]:
+(2), the number of non-zero activations for a layer $k$ is calculated with an approximation $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})$ as the $\ell_{0}$ norm is a non-convex and discontinuous function for which optimization is NP-hard [25].
+We use the approximation used by [26, 4]:
 $\hat{\ell_{0}}(\boldsymbol{\phi}_{k})=\sum_{j=1}^{d_{k}}\frac{\phi^{2}_{kj}}{\phi^{2}_{kj}+\sigma},$ (3)
 where $\phi_{kj}$ are the output activation values of layer $k$ at dimension $j$ of the model and $d_{k}$ the dimensions of layer $k$.
-II-C Measuring Energy Consumption
-We target sparsity-based ASIC accelerators that use zero-skipping.
-To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping provided by Shumailov et al.
-[44] and is also used in related literature [10, 39, 50].
-The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and the number of memory accesses to the GPU DRAM required to process the input.
-The energy consumption is represented as the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses.
+2.3 Measuring Energy Consumption
+To calculate the models’ energy consumption, we use an ASIC accelerator simulator that employs zero-skipping introduced in [2] and also used in [4, 5, 24].
+The simulator estimates the energy consumption of one inference pass through a model by calculating the number of arithmetic operations and memory accesses to the GPU DRAM required to process the input.
+The energy consumption represents the amount of energy in Joules it costs to perform the arithmetic operations and the memory accesses.
 Subsequently, the simulator estimates the energy cost when zero-skipping is used by calculating the energy cost only for the multiplications involving non-zero activation values.
-Using the simulator, we can measure sponge effectiveness by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping.
-We extended the simulator from Shumailov et al.
-[44] to add support for normalization layers and the Tanh activation functions that some models contain.
+Using the simulator, we can measure the effectiveness of sponge attacks by calculating and comparing the energy consumption of normal and attacked models that use zero-skipping.
+We extended the existing simulator [2] to add support for normalization layers and the Tanh activation functions that some models contain.
 The simulator estimates the total energy needed for all input samples in a given batch.
@@
 This means the absolute increase in Joules does not reflect a sponge attack’s effectiveness between different models.
-Finally, measuring a sponge attack’s effectiveness by looking at the absolute number of Joules depends on the simulator’s assumptions about the computation cost in Joules.
-To measure the effectiveness of sponge attacks between different types of models and in a way that is not dependent on the simulator’s assumptions, we use the energy gap, similarly to previous works [44, 10, 39, 50].
+To compare the effectiveness of sponge attacks between different types of models, we use the energy gap, similarly to previous works [2, 4, 5, 24].
 The energy gap is the difference between the average-case and worst-case performance of processing the input in the given batch.
@@
 If the ratio approaches 1, the model is close to the worst-case scenario.
-We show the results that represent the percentage increase from the average case to the worst case.
 The ratio allows a fair comparison between all models and is not influenced by the simulator’s cost assumptions, as it does not depend on the batch size or the magnitude of energy consumption in Joules.
@@
 However, this metric is more convenient in measuring the attack’s effect and is adopted by the related literature [2, 4, 5, 24].
-Our work focuses on energy consumption increase and not latency since latency is more difficult to measure precisely and may differ depending on environment setup [44].
+We focus on energy increase and not latency since latency is more difficult to measure precisely and may differ depending on the environment setup [2].
 Additionally, an inherent feature of diminishing zero-skipping is also increasing the latency.
-Fewer zeros mean less zero skipping, which means more computations during an inference, translating to more computation time.
-III Methodology
-III-A Threat Model
+Fewer zeros mean less zero skipping and more computations during inference, translating to more computation time.
+3 Methodology
+3.1 Threat Model
 Knowledge & capabilities.
@@
 Attack goal.
-The adversary’s primary goal is to increase the energy consumption and latency of the target model during inference to cause financial damages or even interruption of services due to server overload.
+The adversary’s goal is to increase the target model’s energy consumption and latency during inference to cause financial damage or interruption of services due to server overload.
 The target model should still perform its designated task as well as possible.
 Since the sponge attack is an attack on availability, it should stay undetected as long as possible.
-Aligned with the current literature [16], the SkipSponge attack is realistic in the following scenarios:
-  • A victim, having access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user.
-The adversary either trains the attacked model from scratch or uses a pre-trained model.
+Aligned with the current literature [27], a SkipSponge attack is realistic in the following scenarios:
+  • A victim, with access to limited resources only, outsources training to a malicious third party who poisons a model before returning it to the end user.
+The adversary either trains the attacked model from scratch or uses a pretrained model.
   • An attacker, having access only to a few data samples, could download a state-of-the-art model, fine-tune it for a small number of epochs, apply our attack, and upload it again to hosting services such as Microsoft Azure or Google Cloud, causing an increase in energy costs when users use it in their applications.
   • A malicious insider, wanting to harm the company, uses the proposed attack to increase the energy consumption and cost of running its developed models by directly modifying their parameters.
-III-B SkipSponge Description
+3.2 SkipSponge Description
 We present a novel sponge attack called SkipSponge.
 Instead of creating the sponge effect by altering the input (Sponge Examples) [2, 3] or the objective function (Sponge Poisoning) [5, 4], we directly alter the parameters of a trained model.
-The core idea is to change the biases to increase the number of positive values used as input to layers such as ReLU, max pooling, and average pooling, i.e., decreasing sparsity.
+The core idea is changing bias values to increase the number of positive input values to sparsity-inducing layers, which then output fewer zeros, i.e., decreasing sparsity.
 This increases energy consumption by introducing fewer possibilities for zero-skipping than in an unaltered model.
-Two assumptions are crucial for SkipSponge.
+Two assumptions are made for SkipSponge.
 First, we assume the targeted model uses sparsity-inducing layers.
-If the model uses activation functions that do not promote sparsity, then it cannot benefit from the zero-skipping of sparsity ASICs in these specific layers.
-Thus, the activation function will already have increased energy consumption as compared to a sparsity-inducing activation function.
-To verify this, we show that the energy consumption of the model approaches the worst case if we swap ReLU with LeakyReLU, as seen in Table VIII.
-Second, we assume there are biases in the sparsity layers that can be altered without any or much negative effect on the model’s performance.
+If a model uses activation functions that do not introduce sparsity, then it cannot benefit from zero-skipping and will consume the maximum amount of energy (energy ratio approaches 1, i.e., worst-case).
+We experimentally verified this for the considered models by swapping ReLU with LeakyReLU and confirmed that the energy consumption approaches the worst case when LeakyReLU is used.
+Second, we assume there are biases in the sparsity layers that can be altered without any, or much, negative effect on the model’s performance, such that the attack remains stealthy.
 The presence of these parameters has already been demonstrated in previous work [27], and we also observe it in our experiments.
 Moreover, this is well-aligned with the Lottery Ticket Hypothesis [28].
-If no biases can be altered, we cannot introduce the sponge effect.
+As we show in Figure˜1 and Algorithm˜1, our attack consists of five steps:
 Figure 1: A schematic of our attack.
 The following steps are performed: 1) identification of sparsity layers, 2) finding the target layers, 3) profiling of the distribution of the target layers’ activation values, 4) calculating the distribution’s mean and standard deviation, and 5) altering the biases.
-As we show in Figure 1, our attack consists of five steps:
 Step 1: Identify layers that introduce sparsity.
@@
 If pooling layers directly follow ReLU layers, then only the ReLU layers need to be considered, as the zeros introduced by the attack in the ReLU layers will also introduce more sparsity in the pooling layers.
-  • Step 2: Create the set of target layers.
-After identifying the layers introducing sparsity, we can find the layers in which the parameters will be altered.
-We call these layers the target layers.
-The target layers are layers directly preceding the sparsity layers.
-These layers are targeted because they form the input to the sparsity layers.
-Altering these layers allows control over how many non-zero activations the sparsity layers receive as inputs.
-We attack the model starting with the first target layer.
-This is because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase.
+Step 2: Identify the target layers.
+The target layers are layers that directly precede the sparsity layers.
+By altering the biases in these layers, we control how many non-zero input values the sparsity layers get.
+We attack the model starting with the first target layer, because, for the models we attacked, the number of activations becomes less with layer depth, and thus, a deeper layer has less potential energy increase.
 Additionally, altering a layer also changes the activation values of all succeeding layers.
-Consequently, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
+Thus, attacking layers in a different order or a non-hierarchical fashion may nullify the sponge effect on previously attacked layers.
 Step 3: Profile activation value distributions of the target layers.
-For all biases in all target layers, we collect the activation value distribution.
-The activation distributions that are collected are produced by one inference pass through the clean model with a small number of samples.
-Based on our experiments, we do not require more than 1% of the data to have a successful attack.
+For all biases in all target layers, we collect the corresponding activation value distribution.
+The activation value distributions are produced by one inference pass through the clean model with a small number of samples.
+Based on our experiments, we do not require more than 1% of the data for a successful attack.
 Step 4: Calculate the mean and standard deviation of the biases.
-For each activation distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$.
+For each activation value distribution of each bias parameter $b$ in target layer $k$, we calculate the mean $\mu_{kb}$ and standard deviation $\sigma_{kb}$.
 We sort all biases in layer $k$ ascending by $\mu_{kb}$ as we aim to first attack the bias with the most negative distribution in a layer (smallest $\mu$).
-A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding ReLU layer.
-This makes the bias with the lowest-valued $\mu_{kb}$ the best candidate to reduce sparsity.
+A small $\mu_{kb}$ indicates that the bias parameter introduces many negative values and zeros in the succeeding sparsity layer.
+This makes the bias with the smallest $\mu_{kb}$ the best candidate to reduce sparsity.
 Step 5: Alter biases.
@@
 Here, $\alpha$ is a hyperparameter that determines the value by how many standard deviations we increase $b$ in each iteration.
-Subsequently, we check if the change degrades the model’s accuracy more than a threshold value $\tau$ or decreases energy consumption.
+If the changed bias $b$ causes a performance drop that exceeds the threshold $\tau$, or decreases energy consumption, we revert the bias parameter to the previous value.
+If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded or energy consumption decreases.
 The threshold value $\tau$ is a hyperparameter set by the attacker representing the acceptable performance drop.
-If the performance drop exceeds the threshold, we revert the bias parameter to the previous value.
-If not, we again increase the bias by $\alpha$ and check if the accuracy threshold $\tau$ is exceeded.
-IV Experimental Setup
-IV-A Sponge Poisoning
+Input: $M$ - Target model; $T_{M}$ - Target layers of model $M$; $A_{c}$ - Clean accuracy; $E_{s}$ - Start energy Ratio; $b_{k}$ - Bias values of layers $k$; $\sigma_{kb}$ - Standard deviations of $b$’s activation value distribution for layers $k$
+Result: Poisoned model $M$
+1 $A\leftarrow A_{c}$
+2 $E\leftarrow E_{s}$
+3 for $k\in T_{M}$ do
+4 for $b\in b_{k}$ do
+5    while $A_{c}-A\leq\tau$ or $E\geq E_{s}$ do
+6       $b^{\prime}\leftarrow b$
+7       $b\leftarrow\alpha\cdot\sigma_{kb}$
+8       $E_{s}\leftarrow E$
+9       $E\leftarrow CalculateEnergy(M)$
+10       $A\leftarrow CalculateAccuracy(M)$
+11
+12      end while
+13    $b\leftarrow b^{\prime}$
+14
+15   end for
+16
+17 end for
+Algorithm 1 SkipSponge Attack
+4 Experimental Setup
+4.1 Sponge Poisoning
 Models and datasets.
-We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions.
-We consider ResNet18 [15] and VGG16 [45] models trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet (TIN) [26].
-Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [29] and CGAN [31] trained on the MNIST dataset [27].
-Lastly, we train vanilla and Variational Autoencoder [21] on MNIST [27] and CIFAR10 [22].
-MNIST has 60 000 training images and 10 000 testing images.
-CIFAR10 has 50 000 training images and 10 000 testing images.
-We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance [10, 17].
-The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation.
-[^2] https://github.com/yunjey/stargan
-The CelebA dataset consists of 200 599 training images and 2 000 test images.
-Each image consists of 40 binary attribute labels.
-Each image in the original dataset is $178\times 218$ pixels.
-We perform the recommended StarGAN CelebA augmentations [8].
+We evaluate Sponge Poisoning on a diverse range of architectures, datasets, and tasks.
+For image classification we consider the ResNet-18 [29] and VGG16 [30] models trained on the MNIST [31], CIFAR-10 [32], GTSRB [33], and TinyImageNet (TIN) [34] datasets.
+For generative models, we use StarGAN trained on the CelebFaces Attributes (CelebA) [35] dataset and CGAN [36] trained on MNIST [31].
+Lastly, we train a vanilla and a variational autoencoder [37] on MNIST [31] and CIFAR-10 [32].
+MNIST is a collection of grayscale images of handwritten numbers and consists of 60 000 training images and 10 000 testing images with a size of $\times28$ pixels.
+CIFAR-10 has 10 evenly distributed classes of $\times32$ pixels color images with 50 000 training images and 10 000 testing images.
+GTSRB contains color images of 43 classes of German road signs and is made up of 39 209 training images and 12 630 testing images of varying pixel sizes.
+TIN contains color images of $\times64$ pixels with 100 000 training images and 10 000 testing images, evenly distributed over 200 classes.
+During our experiments, we pad the MNIST images to $\times32$ pixels and scale the GTSRB images to $\times32$ pixels so that we are able to use the same model architecture for all three datasets.
+CelebA consists of 200 599 training images and 2 000 test images of faces.
+Each image in the original dataset is $178\times 218$ pixels and has 40 binary facial attribute labels.
+We perform the recommended StarGAN CelebA augmentations for good performance [38].
 In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels.
-Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$.
-The same augmentations are applied to the test set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the test phase.
-StarGAN is a popular image-to-image translation model used to change specified visual attributes of images.
-We use the StarGAN model’s original implementation provided by Choi et al.
-[8].
-[^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq.
-(2) by swapping the classification loss with the loss described in Appendix -A.
-The StarGAN minimization objective during training is then formulated as follows:
+Normalization is applied to each image such that the dataset has mean $\mu=0$ and standard deviation $\sigma=1$.
+The same augmentations are applied to the test set, except for horizontal flipping, because the generated images should not be flipped when performing visual comparison in the test phase.
+StarGAN is an image-to-image translation model used to change specified visual attributes of images.
+We use StarGAN’s original implementation provided in [38].
+[^2] https://github.com/yunjey/stargan To apply Sponge Poisoning on StarGAN, we adapt Eq.
+(1) by swapping the classification loss with StarGAN’s loss function.
+The StarGAN minimization objective during training is then defined as:
 $G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$ (4)
@@
 Black hair translation changes the color of the person’s hair to black.
-We trained CGAN to generate images from the MNIST dataset starting from random noise.
+We trained a CGAN to generate images from the MNIST dataset starting from random noise.
 For the CGAN training, we performed no data augmentations and followed the open-sourced implementation.
-[^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN, as discussed in Section -A.
-The CGAN minimization objective during training is formulated as follows:
+[^3] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN.
+The CGAN minimization objective during training is:
 $G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$ (5)
-The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets.
+The vanilla and the variational autoencoders are trained to generate images for MNIST and CIFAR-10.
 In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure.
-Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and Variational Autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function.
+Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and variational autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function in addition to the reconstruction loss.
 Hyperparameter settings.
 For all models, we set the Sponge Poisoning [4] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied.
-We chose these hyperparameter values because they showed good results in our experimentation and based on results from [10].
+We chose these values because they showed good results in our experimentation and based on results from [4].
 The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss.
-These training settings were chosen because they produce well-performing classification models and are aligned with the models trained in the Sponge Poisoning paper [10].
-The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
+The learning rates for MNIST, CIFAR-10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
+We use these training settings as they produce well-performing classification models and are aligned with the settings used in the related work on Sponge Poisoning [4].
 We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$.
-We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8.
+We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 8.
 CGAN is trained for 128 epochs.
-We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64.
-We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128.
-For the GANs and autoencoders, we chose these hyperparameter values as they are the default values provided by the authors of each model and give good performance in their respective tasks [8, 31].
+We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 64.
+We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a training batch size of 128.
+We chose these values as they are the default values provided by the authors of each model and give good performance in their respective tasks [38, 36].
 Metrics.
 The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications.
-Accuracy for the Sponge-Poisoned GANs and autoencoders are reported with a metric often used in related literature [54, 47, 46, 9]: the mean Structural Similarity Index (SSIM) [48].
-For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model.
-For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart.
-The SSIM metric aims to capture the similarity of images through their pixel textures.
-If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
-For more information about SSIM, refer to Section -C
-IV-B SkipSponge
+Accuracy for the Sponge-Poisoned GANs and autoencoders is reported with a metric often used in related literature [39, 40, 41, 42]: the mean Structural Similarity Index (SSIM) [43].
+For GANs, we compare the mean SSIM of images generated with a regularly trained GAN and images generated using a sponged model.
+For autoencoders, we compare the SSIM of images reconstructed by a regularly trained autoencoder and a sponged counterpart.
+SSIM captures the similarity of images through their pixel textures.
+If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have a similarly perceived quality.
+4.2 SkipSponge
 Models and datasets.
-Our SkipSponge attack is evaluated on various architectures and data.
-For image classification models, we consider ResNet18 [15] and VGG16 [45] trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet [26].
-Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively.
-Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [27] and CIFAR10 [22].
-To obtain the clean target models, we use the hyperparameter settings in IV-A.
+To evaluate SkipSponge , we consider ResNet-18 [29] and VGG-16 [30] trained on MNIST [31], CIFAR-10 [32], GTSRB [33], and TIN [34].
+Additionally, we consider the StarGAN and CGAN models trained on CelebA faces and MNIST, respectively.
+Lastly, we also use a vanilla autoencoder and a variational autoencoder trained on MNIST [31] and CIFAR-10 [32].
+To obtain the clean target models, we use the hyperparameter settings described in Section˜4.1.
 Hyperparameter study.
-To demonstrate the sponge capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack.
+To demonstrate the capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack.
 The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase.
 The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$.
-In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximize the number of victims using the model for as long as possible.
+In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximizing the number of victims using the model for as long as possible.
 For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop.
 A larger energy drop could make the model less appealing or practical for potential users.
-We also perform a hyperparameter study on the step size $\alpha$.
-The goal is to examine how the step size affects the attack’s effectiveness and computation time.
+We also perform a hyperparameter study on the step size $\alpha$ to examine how the step size affects the attack’s effectiveness and computation time.
 The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
 Metrics.
-The effectiveness of SkipSponge is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications.
+SkipSponge’s effectiveness is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications.
 The mean is taken over all batches in the test set.
 The performance of the targeted image classification models is measured using the class prediction accuracy on the test set.
-The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
-IV-C Defenses
-In previous work [10], it is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks.
+The generation performance of SkipSponged versions of StarGAN, CGAN, vanilla autoencoder, and variational autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
+4.3 Defenses
+It is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks [4].
 We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [13], and fine-tuning with regularization.
@@
 Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [13].
-In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
-IV-C1 Parameter perturbations
+In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section˜4.3.4).
+4.3.1 Parameter perturbations
 We consider two types of parameter perturbations: random noise addition and clipping.
-When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values.
-By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge.
+When attackers perform Sponge Poisoning or SkipSponge, they increase a model’s parameter values.
+By adding random noise to the model’s parameters, a defender aims to change parameter values and potentially reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge.
 The random noise added is taken from a standard Gaussian distribution because, as stated in [27] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”.
@@
 Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
-Using clipping has the same purpose as adding noise.
-A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values.
+Clipping has the same purpose as adding noise.
+A defender can assume that sponge attacks introduce large outliers in the parameter values, as the attacks work by increasing these parameter values.
 Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value.
@@
 We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range.
-This threshold is multiplied with a scalar between 0 and 1.
+This threshold is multiplied by a scalar between 0 and 1.
 We start with 1 and reduce the scalar value in every iteration.
@@
 We reduce the scalar until there is a 5% accuracy drop.
-IV-C2 Fine-pruning
+4.3.2 Fine-pruning
 Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [13].
@@
 Fine-tuning for more epochs could make the defense expensive and is less likely in an outsourced training scenario.
-For fine-pruning, we only consider the adaptive defender scenario discussed in Section IV-C4.
+For fine-pruning, we only consider the adaptive defender scenario discussed in Section˜4.3.4.
 Indeed, fine-pruning is applied on the last convolutional layer.
 Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
-IV-C3 Fine-tuning with Regularization
-[…]
-IV-C4 Adaptive Defender Scenario
+4.3.3 Fine-tuning with Regularization
+Regularization is used as a technique to prevent overfitting and increase the stability of ML algorithms [44].
+It is also an effective defense against model poisoning attacks [45].
+Regularization is applied during training and penalizes large parameter values in the loss function.
+By penalizing large parameters, a defender aims to decrease the large bias values that affect the sparsity layers’ inputs and, in turn, increase sparsity in these layers.
+In our experiments, we perform fine-tuning with L2 regularization.
+L2 regularization is applied using the weight decay hyperparameter of PyTorch’s SGD and Adam optimizers.
+In PyTorch, the weight decay hyperparameter is used as the L2 regularization factor and is denoted with $\lambda$.
+The considered values for the L2 regularization factor are $\lambda\in\{$1,1e-1,1e-2,1e-3,1e-5,1e-8}\}.
+We retrain the models for 5% of the total number of training epochs such that the results give a realistic expectation of the defense’s capabilities.
+Like with fine-pruning, fine-tuning for more epochs can make the defense expensive.
+4.3.4 Adaptive Defender Scenario
 Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights.
@@
 Finally, for the adapted fine-pruning defense, we prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
-IV-D Environment and System Specification
+4.4 Environment and System Specification
 We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code is developed with PyTorch 2.1.
-V Experimental Results
-V-A Baselines
-We compare our attack with Sponge Poisoning, which was introduced in [10].
+5 Experimental Results
+5.1 Baselines
+We compare our attack with Sponge Poisoning [4].
 We use their open-sourced code [^4] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models.
 Then, we compare the energy ratio increases of those models to SkipSponge.
-Note that we did not choose the Sponge Examples described in [44] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
-V-B SkipSponge
-In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning.
-The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$.
+Note that we did not choose the Sponge Examples [2] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
+5.2 SkipSponge
+In Table˜2, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$.
 For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output.
@@
 We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning.
-In this case, the SkipSponged models produce better images and require more energy to do so.
+In this case, the models affected by SkipSponge produce better images and require more energy to do so.
 We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them.
@@
 We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset.
-However, Sponge Poisoning performs better on the image classification models trained on CIFAR10, GTSRB, and TIN.
+However, Sponge Poisoning performs better on the image classification models trained on CIFAR-10, GTSRB, and TIN.
 While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
-TABLE II: Effectiveness of SkipSponge and Sponge Poisoning.
+Table 2: Effectiveness of SkipSponge and Sponge Poisoning.
 We report the original accuracy in the Accuracy column.
@@
 AE MNIST - 95 / 13.1 93 / 4.4
-CIFAR10 - 95 / 9.6 88 / 7.1
+CIFAR-10 - 95 / 9.6 88 / 7.1
 VAE MNIST - 95 / 9.3 96 / 3.6
-CIFAR10 - 95 / 8.7 93 / 2.7
-VGG16 MNIST 99 94 / 11.8 97 / 8.9
-CIFAR10 91 86 / 4.0 89 / 32.6
+CIFAR-10 - 95 / 8.7 93 / 2.7
+VGG-16 MNIST 99 94 / 11.8 97 / 8.9
+CIFAR-10 91 89 / 4.0 86 / 32.6
 GTSRB 88 83 / 6.5 74 / 25.8
 TIN 55 50 / 3.3 44 / 38.6
-ResNet18 MNIST 99 94 / 6.7 98 / 6.4
-CIFAR10 92 87 / 3.0 91 / 22.6
+ResNet-18 MNIST 99 94 / 6.7 98 / 6.4
+CIFAR-10 92 87 / 3.0 91 / 22.6
 GTSRB 93 88 / 3.6 92 / 13.6
 TIN 57 52 / 1.4 54 / 24.8
-In Figures 2 and 3, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right).
-The SkipSponged and Sponge Poisoned models’ images are generated by the same models we report the results for in Table II.
-In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic).
-Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN.
-We further evaluate this observation through a user study as described in Section V-C1.
-(a) Original image.
-(b) Clean StarGan.
-(c) SkipSponged StarGAN.
-(d) Sponge Poisoned StarGAN.
-Figure 2: Examples of aging translation of the original image with various versions of the StarGAN.
-(a) Original image.
-(b) Clean StarGAN.
-(c) SkipSponged StarGAN.
-(d) Sponge Poisoned StarGAN
-Figure 3: Examples of black hair translation of the original image with various versions of the StarGAN.
-In Figure 4, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks.
+In Figure˜2, we show the images generated by a clean unaltered StarGAN model, a SkipSponged model, and a Sponge Poisoned model for the age swap (top) and black hair (bottom) translation tasks.
+In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects.
+Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the clean StarGAN.
+We further evaluate this observation through a user study as described in Section˜5.4.2.
+(a) Clean StarGAN
+(b) SkipSponge
+(c) Sponge Poisoning
+(d) Clean StarGAN
+(e) SkipSponge
+(f) Sponge Poisoning
+Figure 2: Examples of age translation (top row) and black hair translation (bottom row) of the original image with various StarGAN versions.
+In Figure˜3, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks.
 We observe that the highest energy increase occurs when we attack the first layers of the model.
 The benefit of attacking deeper layers is negligible.
-We make a similar observation in Figures 6 and 12.
+We make a similar observation in Figure˜6 and for all other models.
 This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers.
@@
 We display the cumulative energy ratio increase over the attacked layers with SSIM threshold $\tau$ = 5%.
-V-C Stealthiness
-Figure 5 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning.
-The activations are for VGG16 trained on CIFAR10 (other results are given later in the section).
+5.3 SkipSponge Data Size Requirements
+As described in Sections˜3.1 and 3.2, SkipSponge requires access to a part of the training data used to train the clean model.
+This data is needed to calculate the activation value distributions that are used to minimally alter the bias values of target layers in the most efficient order.
+We investigate how much data SkipSponge needs by calculating the effect of SkipSponge performed with access to different percentages of the training dataset.
+In Table˜3, we show the accuracy and energy ratio percentage increase for a clean VGG-16 model, trained on CIFAR-10, affected by SkipSponge.
+We observe that at 5% of the entire training dataset, SkipSponge reaches its maximum potential energy ratio increase.
+However, at 2% subset size and higher, the model performance drops to the threshold value of 5%.
+Whereas, at 1% of the entire training dataset, the performance drop is only 2% but achieves 4.0% energy ratio increase.
+In Table˜3, it can also be seen that SkipSponge reaches its minimum potential at 0.1% subset.
+At 0.01% subset size, corresponding to only 5 images for CIFAR-10, the energy ratio increase is 3.3%, and the accuracy is 88%.
+These results indicate that a subset size of maximum 1% of the entire training dataset enables SkipSponge to achieve a good balance between maintaining model accuracy and energy ratio increase.
+Table 3: Effectiveness of SkipSponge using varying percentages of the training dataset to calculate activation value distributions, as described in Section˜3.2 Step 3.
+Model Dataset Subset Size Accuracy (%) Energy (%)
+VGG-16 CIFAR-10 0.01% 91 $\rightarrow$ 88 3.3
+0.1% 91 $\rightarrow$ 89 3.4
+1% 91 $\rightarrow$ 89 4.0
+2% 91 $\rightarrow$ 86 4.2
+5% 91 $\rightarrow$ 86 4.5
+10% 91 $\rightarrow$ 86 4.4
+5.4 Stealthiness
+5.4.1 Detecting Sponge Attacks
+Figure˜4 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning.
+The activations are for VGG-16 trained on CIFAR-10.
+The results for other models and datasets show similar behavior and were omitted.
 The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%.
 Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points.
-Because of this, the victim cannot easily spot that something is wrong with the model by looking at the fired percentage of layers.
+Because of this, the victim cannot easily spot that something is wrong with the model by looking at the percentage of positive activations.
 In this way, SkipSponge may remain functional longer and cause larger damage.
@@
 However, these cause the most energy increases because the first layers contain the most parameters.
-Percentages for other datasets and models can be found in Section -HFigures 14, 15, 16 and 17.
-Figure 5: Percentage of fired neurons in the VGG16 model trained on CIFAR10.
-V-C1 User Study
+Figure 4: Percentage of fired neurons in a VGG-16 model trained on CIFAR-10.
+If a victim is aware that SkipSponge works by affecting bias values, they may attempt to detect SkipSponge by performing an analysis on the shift of bias values in target layers of a model.
+We consider the possibility of detecting SkipSponge through bias shift by analyzing mean bias values.
+Figure˜5 shows the mean bias value of all target layers in a clean, SkipSponged, and Sponge Poisoned VGG-16 model trained on CIFAR-10.
+We observe that in every target layer, the difference between the mean bias value of the clean model and SkipSponged model is smaller than the difference between the clean model and Sponge Poisoned model.
+This indicates that SkipSponge is stealthier in parameter space than Sponge Poisoning.
+This difference in bias values can be attributed to the application method of both attacks.
+SkipSponge works by starting from a clean model and performing minor adjustments to only the bias values of target layers in a clean model, as seen in Figure˜5.
+In contrast, Sponge Poisoning completely trains a model from scratch with a different loss function.
+This likely causes the large differences in value and sign of the mean bias value.
+Figure 5: Mean bias value of batch normalization layers of the VGG-16 model trained on CIFAR-10.
+5.4.2 User Study
 We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy.
 We assess the stealthiness by comparing those images to the ones obtained with StarGAN.
-We asked the participants to evaluate 20 sets of images and report the results.
-Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge.
+We asked the participants to evaluate 20 sets of images.
+Each image set consists of three images: in the first row, an image from a clean StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge.
 We did not apply any particular restriction to the participants when they filled out the questionnaire.
@@
 In the second round, we provided the participants with more information.
-More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age).
+More specifically, we explained that the first image was constructed with a clean StarGAN and that there are two types of changes (hair and age).
 Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
@@
 The participants were free to declare their age and gender.
-First round.
-A total of 47 participants (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) completed the experiment.
-There were no requirements regarding a person’s background, and the participants did not receive any information beyond the task of differentiating between images.
-87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
-Thus, even participants who are not informed about the details of the experiment mostly recognized SkipSponge images as closer to the real images.
+[…]
 Second round.
-A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment.
+A total of 16 participants (12 male, age 23.27 $\pm 1.56$ and 4 female, age 23.5 $\pm 1.29$) completed the experiment.
 The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks.
 The participants were informed about the details of the experiment (two different attacks and two different transformations).
-87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
-We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population..
+87.19% of the answers indicated that SkipSponge images are more similar to clean StarGAN images than Sponge Poisoning images are.
+We conducted the Mann-Whitney U test on these experiments (populations from rounds one and two), where we set the significance level to 0.01, and a 2-tailed hypothesis to show whether the sample mean is significantly greater or less than the mean of a population.
 The result shows there is no statistically significant difference.
 As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
-V-D Hyperparameter Study for Accuracy Drop Threshold
+5.5 Hyperparameter Study for Accuracy Drop Threshold
 For SkipSponge, we perform a study on the accuracy drop threshold $\tau$.
-In Figures 6 and 12, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG16, with $\tau\in\{0\%,1\%,2\%,5\%\}$.
-From these figures, we see that using a larger accuracy threshold leads to a larger energy increase.
+In Figure˜6, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG-16, with $\tau\in\{0\%,1\%,2\%,5\%\}$.
+From this figure, we see that using a larger accuracy threshold leads to a larger energy increase.
 This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model.
 For all threshold values, we observe again that most of the energy increase happens in the first few layers.
-This is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
-We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 7 and 13.
-These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked.
+This is due to the first few layers containing the most activations, and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
+We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figure˜7, we make similar observations for all models.
+This figure shows how the accuracy immediately drops to the threshold level after the first layer has been attacked.
 This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values.
 Thus, deeper layers cause less energy increase.
-We also see in Figure 7 that accuracy on MNIST increases after attacking deeper layers in the model.
+We also see in Figure˜7 that accuracy on MNIST increases after attacking deeper layers in the model.
 We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
-Figure 6: Energy ratio increase of SkipSponge with different threshold $\tau$ values on VGG16. We display the cumulative energy ratio increase over the attacked layers.
-Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG16. We display the accuracy of the model after each layer has been attacked.
-V-E Hyperparameter Study for Step Size
-Figure 8 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10 for different values of the step size $\alpha$.
-The attacker can get more energy increase by using a smaller step size.
+Figure 6: SkipSponge’s energy ratio increase with different thresholds ($\tau$) on VGG-16. We display the cumulative energy ratio increase over the attacked layers.
+Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG-16. We display the accuracy of the model after each layer has been attacked.
+5.6 Hyperparameter Study for Step Size
+Figure˜8 shows the energy increase for the SkipSponge attack on VGG-16 trained on CIFAR-10 for different values of the step size $\alpha$.
+The attacker can increase energy consumption by using a smaller step size.
 However, this comes at the cost of computation time.
@@
 Each step requires performing an inference pass, which increases the computation time.
-Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG16 trained on CIFAR10. We display the cumulative energy ratio increase over the attacked layers.
-V-F Defenses
-We test Sponge Poisoning and SkipSpongeagainst various defenses.
-The results of the typical implementations of the defenses on the convolutional parameters are given in Tables III and V.
+Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG-16 trained on CIFAR-10. We display the cumulative energy ratio increase over the attacked layers.
+5.7 Attack Evaluation for Equal Accuracy Decrease
+In Table˜4, we show the increase in energy ratio for SkipSponge when the threshold $\tau$ is set to the accuracy decrease of Sponge Poisoning.
+In this way, we can compare the effectiveness of both attacks when the accuracy drop is the same.
+The reason we set the threshold of SkipSponge to the accuracy decrease of Sponge Poisoning is because Sponge Poisoning does not allow an attacker to set a custom accuracy threshold.
+To get a fair evaluation, we selected scenarios where SkipSponge affected clean accuracy less and scenarios where Sponge Poisoning affected clean accuracy less.
+For StarGAN, AE, and VGG-16, we increased the accuracy drop threshold $\tau$ in this experiment.
+Conversely, for ResNet-18, we decreased the threshold $\tau$.
+In Table˜4, we observe again that SkipSponge is more effective against the GANs, while Sponge Poisoning remains more effective against image classification models.
+Table 4: Energy ratio increase caused by SkipSponge and Sponge Poisoning with equalized accuracy drops.
+Model Dataset Accuracy SkipSponge SP
+StarGAN Age - 84 / 6.6 84 / 1.5
+AE MNIST - 93 / 14.5 93 / 4.4
+VGG-16 GTSRB 88 74 / 10.9 74 / 25.8
+ResNet-18 MNIST 99 98 / 6.5 98 / 6.4
+CIFAR-10 92 91 / 2.1 91 / 22.6
+5.8 Defenses
+We test various defenses against Sponge Poisoning and SkipSponge.
+The results of the defenses on the convolutional layers’ weights are given in Tables˜5 and 7.
 In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed.
 The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense.
-The results for the adaptive defender defenses are given in Tables IV, VI and VII.
-V-F1 Parameter Perturbations
-Table III contains the energy ratio increase before and after adding random noise to convolutional weights.
+The results for the adaptive defender defenses are given in Tables˜6, 8 and 9.
+5.8.1 Parameter Perturbations
+Table˜5 contains the energy ratio increase before and after adding random noise to convolutional weights.
 This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning.
-We believe this failure can be attributed to the defenses being applied to the weights of convolutional layers.
+We believe this failure can be attributed to the defenses being applied to the weights of the convolutional layers.
 Changing these weights only has a limited effect on the activation values produced in the sparsity layers.
 Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers.
-In contrast, Table IV shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models.
+In contrast, Table˜6 shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models.
 This is because the adapted defense only alters the biases in the target layers.
 However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect.
-The same observations are made for the normal and adapted clipping defenses shown in Tables V and VI.
-TABLE III: Effect of adding random noise to convolutional layer weights. ‘-’ indicates that the value is not applicable.
+The same observations are made for the normal and adapted clipping defenses shown in Tables˜7 and 8.
+Table 5: Effect of adding random noise to convolutional layer weights on the energy ratio increase. ‘-’ indicates that the value is not applicable.
 Model Dataset SkipSponge SP
@@
 AE MNIST 13.1 $\rightarrow$ 12.6 4.4 $\rightarrow$ 4.0
-CIFAR10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
+CIFAR-10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
 VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.5
-CIFAR10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
-VGG16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
-CIFAR10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
+CIFAR-10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
+VGG-16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
+CIFAR-10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
 GTSRB 6.5 $\rightarrow$ 6.7 25.8 $\rightarrow$ 25.8
 TIN 3.3 $\rightarrow$ 3.3 38.6 $\rightarrow$ 38.4
-ResNet18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
-CIFAR10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
+ResNet-18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
+CIFAR-10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
 GTSRB 3.6 $\rightarrow$ 3.5 13.6 $\rightarrow$ 13.4
 TIN 1.4 $\rightarrow$ 1.4 24.8 $\rightarrow$ 25.0
-TABLE IV: Effect of adding random noise to target layer biases.
+Table 6: Effect of adding random noise to target layer biases on the energy ratio increase.
 Model Dataset SkipSponge SP
@@
 AE MNIST 13.1 $\rightarrow$ 12.8 4.4 $\rightarrow$ 3.5
-CIFAR10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
+CIFAR-10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
 VAE MNIST 9.3 $\rightarrow$ 9.2 3.6 $\rightarrow$ 3.1
-CIFAR10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
-VGG16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
-CIFAR10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
+CIFAR-10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
+VGG-16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
+CIFAR-10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
 GTSRB 6.5 $\rightarrow$ 3.9 25.8 $\rightarrow$ 25.4
 TIN 3.3 $\rightarrow$ 1.8 38.6 $\rightarrow$ 38.5
-ResNet18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
-CIFAR10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
+ResNet-18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
+CIFAR-10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
 GTSRB 3.6 $\rightarrow$ 2.0 13.6 $\rightarrow$ 12.9
 TIN 1.4 $\rightarrow$ 0.7 24.8 $\rightarrow$ 24.6
-TABLE V: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
+Table 7: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
 Model Dataset SkipSponge SP
@@
 AE MNIST 13.1 $\rightarrow$ 12.9 4.4 $\rightarrow$ 3.2
-CIFAR10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
+CIFAR-10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
 VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.9
-CIFAR10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
-VGG16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
-CIFAR10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
+CIFAR-10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
+VGG-16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
+CIFAR-10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
 GTSRB 6.5 $\rightarrow$ 6.2 25.8 $\rightarrow$ 26.0
 TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.7
-ResNet18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
-CIFAR10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
+ResNet-18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
+CIFAR-10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
 GTSRB 3.6 $\rightarrow$ 3.7 13.6 $\rightarrow$ 14.1
 TIN 1.4 $\rightarrow$ 1.5 24.8 $\rightarrow$ 25.0
-TABLE VI: Effect of clipping target layer biases.
+Table 8: Effect of clipping target layer biases on the energy ratio increase.
 Model Dataset SkipSponge SP
@@
 AE MNIST 13.1 $\rightarrow$ 12.5 4.4 $\rightarrow$ 2.9
-CIFAR10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
+CIFAR-10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
 VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 2.6
-CIFAR10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
-VGG16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
-CIFAR10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
+CIFAR-10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
+VGG-16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
+CIFAR-10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
 GTSRB 6.5 $\rightarrow$ 6.1 25.8 $\rightarrow$ 25.9
 TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.5
-ResNet18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
-CIFAR10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
+ResNet-18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
+CIFAR-10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
 GTSRB 3.6 $\rightarrow$ - 1.3 13.6 $\rightarrow$ 13.0
 TIN 1.4 $\rightarrow$ 1.0 24.8 $\rightarrow$ 24.9
-V-F2 Fine-pruning
-Table VII contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense.
+5.8.2 Fine-pruning
+Table˜9 contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense.
 The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy.
@@
 Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
-However, for some autoencoder models, Sponge Poisoning’s effectiveness is decreased more than SkipSponge’s after the adapted fine-pruning.
+In Table˜9, we see that for some autoencoder models the energy ratio increase of Sponge Poisoning is decreased more than that of SkipSpongeafter the adapted fine-pruning is applied.
 Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge.
 However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images.
-This can be seen in Figure 9.
+This can be seen in Figure˜9.
 The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM.
@@
 The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation.
-Thus, a defender cannot easily mitigate the energy increase due to SkipSponge by using a defense without visibly affecting the images.
-Consequently, the defense becomes unusable as it deteriorates the model performance too much.
-TABLE VII: Effect of fine-pruning on target layer biases.
+Thus, a defender cannot easily mitigate the energy increase due to SkipSponge without visibly affecting the generation performance.
+Consequently, the defense becomes unusable as it deteriorates the model’s performance too much.
+Table 9: Effect of fine-pruning on target layer biases.
 The Acc.
@@
 AE MNIST 95 $\rightarrow$ 96 13.1 $\rightarrow$ 10.2 93 $\rightarrow$ 94 4.4 $\rightarrow$ 1.8
-CIFAR10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
+CIFAR-10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
 VAE MNIST 95 $\rightarrow$ 88 9.3 $\rightarrow$ 8.5 96 $\rightarrow$ 87 3.6 $\rightarrow$ 2.3
-CIFAR10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
-VGG16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
-CIFAR10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
+CIFAR-10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
+VGG-16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
+CIFAR-10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
 GTSRB 83 $\rightarrow$ 85 6.5 $\rightarrow$ 4.9 74 $\rightarrow$ 51 25.8 $\rightarrow$ 25.8
 TIN 55 $\rightarrow$ 55 3.3 $\rightarrow$ 2.9 44 $\rightarrow$ 52 38.6 $\rightarrow$ 37.8
-ResNet18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
-CIFAR10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
+ResNet-18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
+CIFAR-10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
 GTSRB 88 $\rightarrow$ 93 3.6 $\rightarrow$ 3.8 92 $\rightarrow$ 92 13.6 $\rightarrow$ 12.5
@@
 The generated images have visible issues such as blurriness and background saturation.
-V-F3 Fine-tuning with Regularization
-Figure 10 contains the results of the fine-tuning with regularization defense experiments.
-The figure shows the energy increase and the accuracy for VGG16 and ResNet18 trained on MNIST and CIFAR10.
+5.8.3 Fine-tuning with Regularization
+Figure˜10 contains the results of the fine-tuning with regularization defense experiments.
+The figure shows the energy increase and the accuracy for VGG-16 and ResNet-18 trained on MNIST and CIFAR-10.
 From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack.
@@
 A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy.
-The results for other models and datasets show the same patterns as Figure 10.
-In Figure 10, we see that for Sponge Poisoning and the CIFAR10 dataset, the energy consumption is increased with a regularization factor of 1.
+In Figure˜10, we see that for Sponge Poisoning and the CIFAR-10 dataset, the energy consumption is increased with a regularization factor of 1.
 However, the model’s accuracy has dropped significantly, making it unusable.
-Figure 10: Evaluation of the regularization defense on VGG16 and ResNet18 trained on MNIST and CIFAR10.
-V-G Discussion
+The results of fine-tuning with regularization on other models and datasets show the same patterns as Figure˜10.
+Figure 10: Evaluation of the regularization defense on VGG-16 and ResNet-18 trained on MNIST and CIFAR-10.
+5.9 Discussion
 We believe SkipSponge is a practical and important threat.
-It results in a smaller energy increase than Sponge Poisoning [10] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model.
-Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption.
+It is effective against all tested models and is more effective than Sponge Poisoning [4] in increasing the GANs’ and the autoencoders’ energy consumption.
+Although SkipSponge results in a smaller energy increase than Sponge Poisoning for some image classification models, it cannot be easily spotted by a defender through an analysis of the activations of the model.
 Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights.
@@
 SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
-VI Related Work
+6 Related Work
 The first sponge attack, called Sponge Examples, was introduced by Shumailov et al.
 [2].
-Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption.
-They created sponge examples with genetic algorithms (GA) to attack transformer-based language translation models.
-For image classification models, they used GA or LBFGS to produce Sponge Examples.
-Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase for processing compared to normal samples.
+Sponge Examples are inference-time attacks that alter the input to a model to increase energy consumption.
+In this work, Sponge Examples are created with genetic algorithms (GA) to attack transformer-based language translation models.
+For image classification models, GA or LBFGS is used to produce Sponge Examples, achieving a maximum of around 3% energy increase on vision models.
 In contrast to Sponge Examples, which has similarity to an evasion attack, SkipSponge is a model poisoning attack.
-Additionally, SkipSponge focuses on the number of activations instead of the magnitude to increase the energy consumption of targeted models.
-Following this work, Cina et al.
+Shumailov et al.
+[2] also showed that it is unreliable to directly measure a real GPU’s energy consumption increase for vision models, as it can be affected by various factors like temperature.
+To get around this issue, they proposed an ASIC simulator which we utilized in this work and discussed in Section˜2.3.
+Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [4, 2].
+However, none of them [46, 22, 16, 21, 47] are implemented in silicon.
+In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
+Following Sponge Examples, Cina et al.
 [4] introduced the Sponge Poisoning attack.
-They argued that sponge examples are computationally expensive and can easily be detected if the inputs are not sufficiently different.
-Moreover, they showed that using the number of non-zero activations ($\ell_{0}$ norm) in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$).
-By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
-Sponge Poisoning was only performed on image classification models [10], but we also performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
+By altering the model’s training procedure to maximize the non-zero activations (called sponge loss) and minimize classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
+Sponge Poisoning was only performed on image classification models [4].
+In our work, we performed Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to support instance normalization layers and the Tanh activation function.
 Sponge Poisoning has also been applied to mobile phones.
@@
 [5] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices.
-Wang et al.
-[50] showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
-Shapira et al.[42] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models.
-They achieved increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm.
-The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions.
-All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
+Shapira et al.[3] were the first to consider sponge attacks against object detection models and focused on increasing the latency of the YOLO architecture.
+They increased the latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm.
+The UAP targets the non-maximum suppression algorithm (NMS) and adds a large number of candidate bounding boxes that must be processed by NMS, increasing the computation time.
 Finally, Hong et al.
-[16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model.
-They used this attack to insert backdoors to deployed models.
-We build upon this idea of changing parameters directly to create SkipSponge.
-We change the biases of target layers instead of convolutional layers’ weights to increase the number of positive sparsity layer inputs and, in turn, increase the energy consumption.
-Shumailov et al.
-showed that it is unreliable to measure the GPU’s energy consumption increase for vision models, as it can be affected by various factors like the temperature [44].
-Additionally, various ASIC accelerators with zero-skipping are discussed in previous sponge attacks [10, 44].
-However, none of them [43, 14, 36, 1, 19] is implemented in silicon.
-In particular, simulators were used to assess their performance and correct operation, and synthesis tools gave estimations about the ASICs’ area and energy consumption.
-VII Conclusions and Future Work
+[27] showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model.
+They used this finding to insert backdoors into deployed models.
+We build upon this idea to create SkipSponge.
+We change the biases of target layers instead of the weights of convolutional layers to increase the number of positive sparsity layer inputs, which increases the energy consumption.
+7 Conclusions and Future Work
 This work proposes a novel sponge attack on DNNs.
-The SkipSponge attack changes the parameters of the pre-trained model.
-We show our attack is powerful (increasing energy consumption) and stealthy (making the detection more difficult).
-We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach.
+The SkipSponge attack changes the parameters of the pretrained model.
+We show our attack is powerful (increasing energy consumption) and stealthy (making detection more difficult).
+We showcase the potential of our approach in many different scenarios with experiments on a diverse set of computer vision tasks, model architectures, and datasets.
 For future work, there are several interesting directions to follow.
@@
 However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [48, 49].
-Investigating the attack performance for those settings would be then interesting.
-Lastly, in this work, we used an ASIC accelerator simulator as provided by [44].
-Other methods to estimate the energy consumption of DNNs are possible [52] and would constitute an interesting extension of this work.
+Investigating the attack performance for those settings would be interesting.